### v3

# Batch learning equals online learning in Bayesian supervised learning

Abstract.
Using functoriality of probabilistic morphisms, we prove that sequential and batch Bayesian inversions coincide in supervised learning models with conditionally independent (possibly non-i.i.d.) data [Le2025]. This equivalence holds without domination or discreteness assumptions on sampling operators. We derive a recursive formula for posterior predictive distributions, which reduces to the Kalman filter in Gaussian process regression. For Polish label spaces $\mathcal{Y}$} and arbitrary input sets $\mathcal{X}$}, we characterize probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ via projective systems, generalizing Orbanz [Orbanz2011]. We revisit MacEachern’s dependent Dirichlet processes (DDP) [MacEachern2000] using copula-based constructions [BJQ2012] and show how to compute posterior predictive distributions in universal Bayesian supervised models with DDP priors.
Key words and phrases:
probabilistic morphism, conditional independence, sequential Bayesian inversion, recursive posterior predictive distribution, Bayesian supervised learning, Gaussian process regression, Dependent Dirichlet process
2020 Mathematics Subject Classification:
Primary: 62C10, Secondary: 62G05, 62G08
1. Introduction
For a measurable space $\mathcal{X}$}, we denote by $\Sigma_{\mathcal{X}}$ the $\sigma$-algebra of $\mathcal{X}$}, and by ${\mathcal{P}}({\mathcal{X}})$ the space of all probability measures on $\mathcal{X}$}. If $({\mathcal{X}},\tau)$ is a topological space, we consider the Borel $\sigma$-algebra ${\mathcal{B}}(\tau)$, denoted also by ${\mathcal{B}}({\mathcal{X}})$, on $\mathcal{X}$}, unless otherwise stated.
In [Le2025] we considered the following problem.
Problem 1.1 (Supervised Bayesian Inference (SBI) Problem).
Let $\mathcal{X}$} be an input space and $\mathcal{Y}$} a measurable label space. Given training data $S_{n}:=\big((x_{1},y_{1}),\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$ and new test data $T_{m}:=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$, estimate the predictive probability measure ${\mathcal{P}}_{T_{m}|S_{n}}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ that governs the joint distribution of the $m$-tuple $\big(y_{1}^{\prime},\ldots,y_{m}^{\prime}\big)\in{\mathcal{Y}}^{m}$ where $y_{i}^{\prime}$ is the label of $t_{i}$.
If $\mathcal{X}$} consists of a single point, Problem SBI is equivalent to the fundamental problem of probability estimation in classical mathematical statistics. If $m=1$ and ${\mathcal{Y}}={\mathbb{R}}^{n}$, under the assumption that the distribution of the label $y$ of $x$ is governed by a corrupted measurement of $f(x)$ for some unknown function $f:{\mathcal{X}}\to{\mathbb{R}}^{n}$, Problem SBI is the regression problem in classical statistics.
In [Le2025], utilizing a categorical approach and stochastic processes taking values in ${\mathcal{P}}({\mathcal{Y}})$ with index set $\mathcal{X}$}, we proposed a Bayesian solution of Problem 1.1 encompassing classical solutions of probability and regression estimation problems that use Bayesian inversions. Our Bayesian modeling (Definition 2.1) of Problem 1.1 works under the assumption of conditionally independent (possibly not identically distributed) data $y\in{\mathcal{Y}}$, which encompasses the classical Bayesian modeling of conditionally i.i.d. data $y\in{\mathcal{Y}}$, assuming $({\mathcal{X}})=1$, see also Remark 2.2(2+3). The classical Bayesian modeling is based on de Finetti’s theorem on exchangeable data and its generalizations.
In this paper, we study posterior distributions, posterior predictive distributions, and universal priors in Bayesian supervised learning. In particular, we prove that batch learning equals online learning in Bayesian supervised learning (Theorems 3.1, 4.1, 4.3).
The question of whether batch learning equals online learning in Bayesian learning has a notable history and importance in mathematical statistics and machine learning. The formal study of updating statistical conclusions one observation at a time (online learning) is known as sequential analysis. The mathematical groundwork for sequential analysis was laid by Abraham Wald [Wald1947]. The explicit formalization of sequential updating in a Bayesian context for estimation problems came to prominence with the development of state-space models and filtering theory [SS2023]. Rudolf E. Kálmán is arguably the most important figure in the practical application of this principle. The Kalman Filter, introduced around 1960, is a perfect example of online Bayesian learning. It uses the posterior from the previous time step as the prior for the current time step to recursively estimate the state of a linear dynamic system. While the Kalman filter is a specific algorithm for Gaussian models, its conceptual basis is precisely the equivalence of batch and online updating. The broader theoretical treatment of this idea in Bayesian statistics is often attributed to Dennis V. Lindley and Adrian F.M. Smith. Their work in the 1970s on Bayesian hierarchical models and the structure of Bayesian inference helped formalize and popularize these recursive computational structures. For instance, their 1972 paper [LS1972] is a landmark in this area. For Bayesian models with conjugate priors, the online approach is simply a recursive way of performing the same computation as the batch approach, breaking it down into smaller, manageable steps without any loss of information or change in the final inference. To the best of our knowledge, until now, the most general available theorem stating that batch Bayesian learning equals online Bayesian learning relies on the assumption of the classical Bayes’ theorem, which assumes either discrete data or sampling operators that are dominated Markov kernels, see, e.g., [SS2023, §3.3].
Our approach is based on properties of compositions of Markov kernels and their graphs, called functoriality of probabilistic morphisms, which we studied in [Le2025] and our characterization of Bayesian inversions as solutions of operator equation for the graph of a Markov kernel (2.8). Prior results on recursive Bayesian updating require additional structure such as dominated kernels or conjugate priors [SS2023]. In contrast, we establish the equivalence of batch and online learning using only conditional independence, functoriality of probabilistic morphisms, and the operator equation characterization of Bayesian inversion developed in [Le2025], combined with the functorial structure of sequential inversion established in Theorem 3.1 of the present work. As a result, we can prove the equivalence without the i.i.d. assumption, holding for the more general case of conditionally independent (possibly not identically distributed) data that our model (Definition 2.1) addresses.
The equivalence is vital because online learning offers advantages such as higher computational efficiency and adaptability, see, e.g., Example 4.4.
Our paper is organized as follows. In Section 2, we recall the concept of probabilistic morphisms, their useful properties, and Bayesian learning models for supervised learning (Definition 2.1). In Section 3, using categorical properties of probabilistic morphisms, we prove that sequential Bayesian inversions in Bayesian supervised learning models is the same as batch Bayesian inversion (Theorem 3.1). Using a projective system, we also derive a formula for Bayesian inversions of a universal supervised learning model $({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ if $\mathcal{X}$} is a finite set (Theorem 3.4). We illustrate Theorems 3.1 and 3.4 by computing the posterior distributions of Dirichlet processes (Example 3.6). In Section 4, we prove recursive formulas for posterior predictive distributions (Theorems 4.1, 4.3) and illustrate Theorem 4.3 with Gaussian process regressions (Example 4.4). Section 5, assuming that $\mathcal{Y}$} is a Polish space and $\mathcal{X}$} is an arbitrary set, characterizes probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ via a projective system, defined by finite subsets in $\mathcal{X}$} and a countable generating algebra of the $\sigma$-algebra of $\mathcal{Y}$} (Theorem 5.5). In Section 6, we illustrate Theorem 5.5 with MacEachern’s Dependent Dirichlet Processes (DDP) priors (Theorem 6.1) and indicate how to compute posterior predictive distributions of universal Bayesian supervised learning models with DDP priors. In the last Section 7, we discuss our results and the concept of predictive consistency in Bayesian supervised learning.
2. Preliminaries
- For a measurable space $\mathcal{X}$}, we denote by $\Sigma_{w}$ the smallest $\sigma$-algebra on ${\mathcal{P}}({\mathcal{X}})$ such that for any $A\in\Sigma_{\mathcal{X}}$ the function $e_{A}:{\mathcal{P}}({\mathcal{X}})\to{\mathbb{R}},\mu\mapsto\mu(A),$ is measurable. In our paper, we always consider ${\mathcal{P}}({\mathcal{X}})$ as a measurable space with the $\sigma$-algebra $\Sigma_{w}$, unless otherwise stated.
- For a measurable space $\mathcal{X}$}, we denote by ${\mathcal{F}}_{b}({\mathcal{X}})$ and ${\mathcal{F}}_{s}({\mathcal{X}})$ the space of measurable bounded functions and the space of all step functions on $\mathcal{X}$}, respectively.
- A Markov kernel $T:{\mathcal{X}}\times\Sigma_{\mathcal{Y}}\to[0,1]$ is uniquely defined by the measurable map $\overline{T}:{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})$ such that $\overline{T}(x)(A)=T(x,A)$ for all $x\in{\mathcal{X}},A\in\Sigma_{\mathcal{Y}}$. We shall also use notations $T(A|x):=T(x,A)$ and $\overline{T}(A|x):=\overline{T}(x)(A)$.
- A probabilistic morphism $T:{\mathcal{X}}\leadsto{\mathcal{Y}}$ is an arrow assigned to a measurable mapping, denoted by $\overline{T}$, from $\mathcal{X}$} to ${\mathcal{P}}({\mathcal{Y}})$. We say that $T$ is generated by $\overline{T}$. For a measurable mapping $T:{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})$ we denote by $\underline{T}:{\mathcal{X}}\leadsto{\mathcal{Y}}$ the generated probabilistic morphism.
- For probabilistic morphisms $T_{{\mathcal{Y}}|{\mathcal{X}}}:{\mathcal{X}}\leadsto{\mathcal{Y}}$ and $T_{{\mathcal{Z}}|{\mathcal{Y}}}:{\mathcal{Y}}\leadsto{\mathcal{Z}}$ their composition is the probabilistic morphism
$T_{{\mathcal{Z}}|{\mathcal{X}}}:=T_{{\mathcal{Z}}|{\mathcal{Y}}}\circ T_{{\mathcal{Y}}|{\mathcal{X}}}:{\mathcal{X}}\leadsto{\mathcal{Z}}$
$(T_{{\mathcal{Z}}|{\mathcal{Y}}}\circ T_{{\mathcal{Y}}|{\mathcal{X}}})(x,C):=\int_{\mathcal{Y}}T_{{\mathcal{Z}}|{\mathcal{Y}}}(y,C)T_{{\mathcal{Y}}|{\mathcal{X}}}(dy|x)$
for $x\in{\mathcal{X}}$ and $C\in\Sigma_{\mathcal{Z}}$. It is well-known that the composition is associative.
- We denote by ${\bf Meas}({\mathcal{X}},{\mathcal{Y}})$ the set of all measurable mappings from a measurable space $\mathcal{X}$} to a measurable space $\mathcal{Y}$}, and by ${\bf Probm}({\mathcal{X}},{\mathcal{Y}})$ the set of all probabilistic morphisms from $\mathcal{X}$} to $\mathcal{Y}$}. We regard ${\bf Meas}({\mathcal{X}},{\mathcal{Y}})$ as a subset of ${\bf Probm}({\mathcal{X}},{\mathcal{Y}})$, identifying ${\mathcal{Y}}\ni y$ with the Dirac measure $\delta_{y}\in{\mathcal{P}}({\mathcal{Y}})$. This is possible, since the Dirac map ${\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}),y\mapsto\delta_{y}$, is measurable [Lawvere62],[Giry82, Theorem 1].
- For $T_{i}\in{\bf Probm}({\mathcal{X}}_{i},{\mathcal{X}}_{i+1})$, $i=1,2$, we have [Chentsov72, Lemma 5.5]
(2.1) $\overline{T_{2}\circ T_{1}}=(T_{2})_{*}\circ\overline{T_{1}}.$
- We denote by ${\mathcal{Y}}^{\mathcal{X}}$ the set of all mappings from $\mathcal{X}$} to $\mathcal{Y}$}. If $\mathcal{Y}$} be measurable space, then ${\mathcal{Y}}^{\mathcal{X}}$ is a measurable space with the cylindrical $\sigma$-algebra denoted by $\Sigma_{cyl}({\mathcal{Y}}^{\mathcal{X}})$.
- For any $\mathcal{X}$} we denote by $\operatorname{Id}_{\mathcal{X}}$ the identity map on $\mathcal{X}$}. For a product space ${\mathcal{X}}\times{\mathcal{Y}}$ we denote by $\Pi_{\mathcal{X}}$ the canonical projection to the factor $\mathcal{X}$}.
- For any $T\in{\bf Probm}({\mathcal{X}},{\mathcal{Y}})$ the linear mapping
$P_{*}T:{\mathcal{P}}({\mathcal{X}})\to{\mathcal{P}}({\mathcal{Y}}),$
(2.2) $P_{*}T(\mu)(B):=\int_{\mathcal{X}}\overline{T}(B|x)\,d\mu(x),\,\mu\in{\mathcal{P}}({\mathcal{X}}),\,B\in\Sigma_{\mathcal{Y}},$
is measurable [Lawvere62], [Giry82, Theorem 1]. If $\mathcal{X}$} and $\mathcal{Y}$} are Polish spaces, then ${\mathcal{P}}({\mathcal{X}})$ and ${\mathcal{P}}({\mathcal{Y}})$ are Polish spaces endowed with the weak*-topology $\tau_{w}$, and their $\sigma$-algebra $\Sigma_{w}$ is the Borel $\sigma$-algebra ${\mathcal{B}}(\tau_{w})$. Furthermore, $P_{*}T$ is a $(\tau_{w},\tau_{w})$-continuous map [Giry82, Theorem 1]. Moreover, for any $T_{1}\in{\bf Probm}({\mathcal{X}}_{1},{\mathcal{X}}_{2})$, $T_{2}\in{\bf Probm}({\mathcal{X}}_{2},{\mathcal{X}}_{3})$ we have [Lawvere62], [Chentsov72, Lemma 5.10],[Giry82, Theorem 1], [JLT21, Proposition 5]
(2.3) $P_{*}(T_{2}\circ T_{2})=P_{*}T_{2}\circ P_{*}T_{1}.$
We also use the abbreviation $T_{*}$ for $P_{*}T$.
- For any $k\in{\mathbb{N}}^{+}$ the multiplication mapping
(2.4) $\mathfrak{m}^{k}:\prod_{i=1}^{k}\big({\mathcal{P}}({\mathcal{X}}_{1}),\Sigma_{w}\big)\to\Big({\mathcal{P}}\big(\prod_{i=1}^{k}{\mathcal{X}}_{i}\big),\Sigma_{w}\Big),\,(\mu_{1},\ldots,\mu_{k})\mapsto\otimes_{i=1}^{k}\mu_{i}$
is measurable [Le2025, Proposition 2.1 (1)].
- For a probabilistic morphism $T:{\mathcal{X}}\leadsto{\mathcal{Y}}$ the graph $\Gamma_{T}:{\mathcal{X}}\leadsto{\mathcal{X}}\times{\mathcal{Y}}$ of $T$ is defined as follows:
$\overline{\Gamma_{T}}(x):=\mathfrak{m}^{2}(\overline{\operatorname{Id}_{\mathcal{X}}},\overline{T}).$
Note that $\overline{\operatorname{Id}_{\mathcal{X}}}=\delta\circ\operatorname{Id}_{\mathcal{X}}$, where
$\delta:{\mathcal{X}}\to{\mathcal{P}}({\mathcal{X}})$
is the measurable map assigning $x\in{\mathcal{X}}$ to the Dirac measure $\delta_{x}$ concentrated at $x$. For any $\mu\in{\mathcal{P}}({\mathcal{X}})$, $A\in\Sigma_{\mathcal{X}}$, $B\in\Sigma_{\mathcal{Y}}$ we have [Le2025, (2.13)]
(2.5) $(\Gamma_{T})_{*}\mu(A\times B)=\int_{\mathcal{X}}\overline{\Gamma_{T}}(x)(A\times B)d\mu(x)=\int_{A}\overline{T}(B|x)\,d\mu(x).$
By [Le2025, Lemma 2.10 (1), (2.15)], we have the following decomposition:
(2.6) $T=\Pi_{\mathcal{Y}}\circ\Gamma_{T}.$
In [Le2025, Lemma 2.10(2)] Lê proved the following formula for the graph of a composition of probabilistic morphisms $p_{1}:{\mathcal{X}}\leadsto{\mathcal{Y}}$ and $p_{2}:{\mathcal{Y}}\leadsto{\mathcal{Z}}$
(2.7) $\Gamma_{p_{2}\circ p_{1}}=(\operatorname{Id}_{\mathcal{X}}\times p_{2})\circ\Gamma_{p_{1}}.$
- A Bayesian statistical model is a quadruple $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{X}})$, where $(\Theta,\mu_{\Theta})$ is a probability space, and ${\mathbf{p}}\in{\bf Meas}\big(\Theta,{\mathcal{P}}({\mathcal{X}})\big)$. The predictive distribution $\mu_{\mathcal{X}}\in{\mathcal{P}}({\mathcal{X}})$ of a Bayesian statistical model $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{X}})$ is defined as the prior marginal probability of $x$, i.e., $\mu_{\mathcal{X}}:=(\Pi_{\mathcal{X}})_{*}\mu$, where $\mu:=(\Gamma_{\underline{{\mathbf{p}}}})_{*}\mu_{\Theta}\in{\mathcal{P}}(\Theta\times{\mathcal{X}})$ is the joint distribution of $\theta\in\Theta$ and $x\in{\mathcal{X}}$ whose regular conditional probability measure with respect to the projection $\Pi_{\Theta}:\Theta\times{\mathcal{X}}\to\Theta$ is ${\mathbf{p}}:\Theta\to{\mathcal{P}}({\mathcal{X}})$. A Bayesian inversion ${\mathbf{q}}:={\mathbf{q}}(\cdot\|{\mathbf{p}},\mu_{\Theta})\in{\bf Meas}\big({\mathcal{X}},{\mathcal{P}}(\Theta)\big)$ of a Markov kernel ${\mathbf{p}}\in{\bf Meas}\big(\Theta,{\mathcal{P}}({\mathcal{X}})\big)$ relative to $\mu_{\Theta}$ is a Markov kernel such that
(2.8) $(\sigma_{{\mathcal{X}},\Theta})_{*}(\Gamma_{\underline{{\mathbf{q}}}})_{*}\mu_{\mathcal{X}}=(\Gamma_{\underline{{\mathbf{p}}}})_{*}\mu_{\Theta},$
where $\sigma_{{\mathcal{X}},\Theta}:{\mathcal{X}}\times\Theta\to\Theta\times{\mathcal{X}}$ is defined by $(x,\theta)\mapsto(\theta,x)$.
We also write ${\mathbf{q}}(\cdot\|{\mathbf{p}},\mu_{\Theta})$ as ${\mathbf{q}}(\cdot\|\mu_{\Theta})$ if $\mathbf{p}$} is fixed and no confusion can occur.
For $X_{m}:=(x_{1},\ldots,x_{m})\in{\mathcal{X}}^{m}$, we denote by
$E_{X_{m}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{m},h\mapsto\big(h(x_{1}),\ldots,h(x_{m}))\in{\mathcal{P}}({\mathcal{Y}})^{m}$
the evaluation mapping.
For $S_{n}=\big((x_{1},y_{1}),\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, we denote by $\Pi_{\mathcal{X}}(S_{n})$ the ${\mathcal{X}}^{n}$-component of $S_{n}$, namely $\Pi_{\mathcal{X}}(S_{n})=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$. Similarly, $\Pi_{\mathcal{Y}}(S_{n})=(y_{1},\ldots,y_{n})\in{\mathcal{Y}}^{n}$. Now we recall the solution of the problem SBI in [Le2025, Definition 3.2].
Definition 2.1.
A Bayesian learning model for the supervised inference problem SBI consists of a quadruple $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$, where $\mu_{\Theta}\in{\mathcal{P}}(\Theta)$ and ${\mathbf{p}}:\Theta\to{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ is a measurable mapping.
(1) For any $X_{m}=(x_{1},\ldots,x_{m})\in{\mathcal{X}}^{m}$, the Bayesian statistical model $(\Theta,\mu_{\Theta},\mathfrak{m}^{m}\circ E_{X_{m}}\circ{\mathbf{p}},{\mathcal{Y}}^{m})$ parameterizes sampling distributions of $Y_{m}=(y_{1},\ldots,y_{m})\in{\mathcal{Y}}^{m}$, where $y_{i}$ is a label of $x_{i}$, with the sampling operator ${\mathbf{p}}_{X_{m}}:=\mathfrak{m}^{m}\circ E_{X_{m}}\circ{\mathbf{p}}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{m})$.
(2) For a training sample $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, the posterior distribution $\mu_{\Theta|S_{n}}\in{\mathcal{P}}(\Theta)$ after seeing $S_{n}$ is the value ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}\big(\Pi_{\mathcal{Y}}(S_{n})\big)$ of a Bayesian inversion ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}:{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$ of the Markov kernel ${\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{n})$ relative to $\mu_{\Theta}$.
(3) For $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$, the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ of the $m$-tuple $(y^{\prime}_{1},\ldots,y^{\prime}_{m})$ where $y^{\prime}_{i}$ is the label of $t_{i}$, given a training data set $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, is defined as the predictive distribution of the Bayesian statistical model $(\Theta,\mu_{\Theta|S_{n}},{\mathbf{p}}_{T_{m}},{\mathcal{Y}}^{m})$, i.e.,
(2.9) ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}:=(\underline{{\mathbf{p}}_{T_{m}}})_{*}\mu_{\Theta|S_{n}}\in{\mathcal{P}}({\mathcal{Y}}^{m}).$
(4) The aim of a learner is to estimate and approximate the value of the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}$.
Remark 2.2.
(1) A Bayesian inversion ${\mathbf{q}}_{X_{n}}:{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$ of the Markov kernel ${\mathbf{p}}_{X_{n}}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{n})$ relative to $\mu_{\Theta}$ is defined $(\underline{{\mathbf{p}}_{X_{n}}})_{*}\mu_{\Theta}$-a.s. uniquely in the weak sense.  [^1] See [Le2025, Theorem 2.9], for the precise definition. In this paper, we need only Formula (2.8) for a characterization of a Bayesian inversion. In Theorem 4.1 below, we shall show that, given inputs $X_{n}=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$, for each $T_{m}\in{\mathcal{X}}^{m}$, the posterior predictive distribution
(2.10) ${\mathcal{P}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m}),\,(y_{1},\ldots,y_{n})\mapsto{\mathcal{P}}_{T_{m}|\big((x_{1},y_{1}),\ldots(x_{n},y_{n})\big),\mu_{\Theta}},$
is a regular conditional probability measure of the joint distribution
$(\underline{\mathfrak{m}^{2}({\mathbf{p}}_{T_{m}},{\mathbf{p}}_{X_{n}})}_{*}\mu_{\Theta}\in{\mathcal{P}}({\mathcal{Y}}^{m}\times{\mathcal{Y}}^{n})$. Hence, ${\mathcal{P}}^{n}_{m}$ is defined $(\underline{{\mathbf{p}}_{X_{n}}})_{*}\mu_{\Theta}$-a.s. uniquely in the weak sense, independently from the choice of a Bayesian inversion ${\mathbf{q}}_{X_{n}}:{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$.
(2) If $({\mathcal{X}})=1$, our Bayesian learning model is a classical Bayesian statistical model $\big(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{P}}({\mathcal{Y}})\big)$ for Bayesian inference under the assumption of conditionally i.i.d. data $y\in{\mathcal{Y}}$.
(3) We showed in [Le2025] that classical Bayesian regression learning is a particular case of Bayesian supervised learning in the sense of Definition 2.1, see Section 4 and Theorem 4.3.
(4) It is easy to verify, e.g., by using Equation (3.4) in the next Section, that our solution of Problem SBI satisfies the following consistency property. For any $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$, $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, and $t_{m+1}\in{\mathcal{X}}$ we have
(2.11) ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}=(\Pi_{{\mathcal{Y}}^{m}})_{*}{\mathcal{P}}_{T_{m},t_{m+1}|S_{n},\mu_{\Theta}}.$
In particular, to solve Problem SBI for all $m\in{\mathbb{N}}^{+}$, it suffices to solve the problem for the case $m=\infty$.
In [Le2025, Proposition 3.4] we showed that the quadruple $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\\ \operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$, where $\mu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$, is a universal Bayesian learning model in the sense of Definition 2.1 for solving the problem SBI.
3. Bayesian inversions in Bayesian supervised learning models
In this Section, we prove two theorems (Theorems 3.1 and 3.4) for computing Bayesian inversions in supervised learning model. We illustrate these theorems with Example 3.6.
Theorem 3.1 (Online formula for Bayesian inversion).
Let $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ be a Bayesian model for supervised learning. Let $S_{n}=\big((x_{1},y_{1}),\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$ and $S_{n-1}=\big((x_{1},y_{1}),\ldots,(x_{n-1},y_{n-1})\big)$. Then a Bayesian inversion ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}(\cdot\|\mu_{\Theta}):{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$ of the Markov kernel ${\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{n})$ relative to $\mu_{\Theta}$ can be found by the following formula:
(3.1) ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}(y_{n},\ldots,y_{1}\|\mu_{\Theta}):={\mathbf{q}}_{x_{n}}(y_{n}\|{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}(y_{n-1},\ldots,y_{1})\|\mu_{\Theta}).$
We abbreviate ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}(\cdot\|\mu_{\Theta})$ as ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}$. For the proof of Theorem 3.1 we need Lemma 3.2 below stating that $\underline{{\mathbf{p}}_{x_{n}}}\circ\underline{{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}}:{\mathcal{Y}}^{n-1}\leadsto{\mathcal{Y}}$ is a regular conditional probability measure for $\big(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}}\big)_{*}\mu_{\Theta}\in{\mathcal{P}}({\mathcal{Y}}^{n})$ with respect to the projection $\Pi_{{\mathcal{Y}}^{n-1}}:{\mathcal{Y}}^{n}\to{\mathcal{Y}}^{n-1}$.
[…]
Lemma 3.2.
Assume the condition of Theorem 3.1. Then we have
(3.2) $\big(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}}\big)_{*}\mu_{\Theta}=(\Pi_{{\mathcal{Y}}^{n-1}})_{*}\big(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}}\big)_{*}\mu_{\Theta},$
(3.3) $\big(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}}\big)_{*}\mu_{\Theta}=\big(\Gamma_{\underline{{\mathbf{p}}_{x_{n}}}\circ\underline{{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}}}\big)_{*}\big(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}}\big)_{*}\mu_{\Theta}.$
Proof of Lemma 3.2.
Lemma 3.2 is a particular case of Proposition 3.3 below. ∎
Proposition 3.3.
Assume that $(\Theta,\mu_{\Theta})$ is a probability space, ${\mathcal{X}},{\mathcal{Y}}$ are measurable spaces. Let ${\mathbf{p}}_{\mathcal{X}}\in{\bf Meas}\big(\Theta,{\mathcal{P}}({\mathcal{X}})\big)$, ${\mathbf{p}}_{\mathcal{Y}}\in{\bf Meas}\big(\Theta,{\mathcal{P}}({\mathcal{Y}})\big)$. If ${\mathbf{q}}_{\mathcal{Y}}\in{\bf Meas}({\mathcal{Y}},{\mathcal{P}}(\Theta))$ is a Bayesian inversion of ${\mathbf{p}}_{\mathcal{Y}}$ relative to $\mu_{\Theta}$, then we have
(3.4) $\big(\underline{{\mathbf{p}}_{{\mathcal{Y}}}}\big)_{*}\mu_{\Theta}=(\Pi_{{\mathcal{Y}}})_{*}\big(\underline{\mathfrak{m}^{2}({\mathbf{p}}_{\mathcal{Y}},{\mathbf{p}}_{\mathcal{X}})}\big)_{*}\mu_{\Theta},$
(3.5) $\big(\underline{\mathfrak{m}^{2}({\mathbf{p}}_{\mathcal{Y}},{\mathbf{p}}_{\mathcal{X}})}\big)_{*}\mu_{\Theta}=\big(\Gamma_{\underline{{\mathbf{p}}_{\mathcal{X}}}\circ\underline{{\mathbf{q}}_{\mathcal{Y}}}}\big)_{*}\big(\underline{{\mathbf{p}}_{\mathcal{Y}}}\big)_{*}\mu_{\Theta}.$
[…]
Proof of Proposition 3.3.
(1) To prove (3.4), it suffices to show that
(3.6) ${\mathbf{p}}_{\mathcal{Y}}=\overline{\Pi_{\mathcal{Y}}\circ\underline{\mathfrak{m}^{2}({\mathbf{p}}_{\mathcal{Y}},{\mathbf{p}}_{\mathcal{X}})}}.$
Using Composition Formula (2.1), we conclude that Equation (3.6) is equivalent to the following:
(3.7) ${\mathbf{p}}_{\mathcal{Y}}(\theta)=(\Pi_{{\mathcal{Y}}})_{*}\mathfrak{m}^{2}({\mathbf{p}}_{\mathcal{Y}},{\mathbf{p}}_{\mathcal{X}})(\theta)\;\forall\theta\in\Theta.$
Equation (3.7) holds since $\mathfrak{m}^{2}({\mathbf{p}}_{\mathcal{Y}},{\mathbf{p}}_{\mathcal{X}})(\theta)={\mathbf{p}}_{\mathcal{Y}}(\theta)\otimes{\mathbf{p}}_{\mathcal{X}}(\theta)$.
(2) Using Formula (2.7) for the graph of a composition of probabilistic morphisms, we obtain
$\big(\Gamma_{\underline{{\mathbf{p}}_{{\mathcal{X}}}}\circ\underline{{\mathbf{q}}_{\mathcal{Y}}}}\big)_{*}\big(\underline{{\mathbf{p}}_{\mathcal{Y}}}\big)_{*}\mu_{\Theta}$
(3.8) $=\big(\operatorname{Id}_{{\mathcal{Y}}}\times\underline{{\mathbf{p}}_{{\mathcal{X}}}}\big)_{*}\big(\Gamma_{\underline{{\mathbf{q}}_{\mathcal{Y}}}})_{*}\big(\underline{{\mathbf{p}}_{\mathcal{Y}}}\big)_{*}\mu_{\Theta}.$
Taking into account that ${\mathbf{q}}_{\mathcal{Y}}:{\mathcal{Y}}\to{\mathcal{P}}(\Theta)$ is a Bayesian inversion of ${\mathbf{p}}_{\mathcal{Y}}:\Theta\to{\mathcal{P}}({\mathcal{Y}})$ relative to $\mu_{\Theta}$, we obtain Equation (3.5) from (3.8):
$\big(\Gamma_{\underline{{\mathbf{p}}_{\mathcal{X}}}\circ\underline{{\mathbf{q}}_{\mathcal{Y}}}}\big)_{*}\big(\underline{{\mathbf{p}}_{\mathcal{Y}}}\big)_{*}\mu_{\Theta}=\big(\operatorname{Id}_{{\mathcal{Y}}}\times\underline{{\mathbf{p}}_{\mathcal{X}}}\big)_{*}(\sigma_{\Theta,{\mathcal{Y}}})_{*}\big(\Gamma_{\underline{{\mathbf{p}}_{\mathcal{Y}}}}\big)_{*}\mu_{\Theta}$
$=\big(\underline{\mathfrak{m}^{2}({\mathbf{p}}_{\mathcal{Y}},{\mathbf{p}}_{\mathcal{X}})}\big)_{*}\mu_{\Theta}.$
Proof of Theorem 3.1.
By Decomposition Formula (2.6), we have
$(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}})_{*}\mu_{\Theta}=(\Pi_{\mathcal{Y}})_{*}(\Gamma_{\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}}})_{*}\mu_{\Theta}.$
Hence, to prove that ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}:{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$ defined by (3.1) is a Bayesian inversion of ${\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{n})$, it suffices to show that
$(\sigma_{\Theta,{\mathcal{Y}}^{n}})_{*}(\Gamma_{\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}}})_{*}\mu_{\Theta}=(\Gamma_{\underline{{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}}})_{*}({\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})})_{*}\mu_{\Theta}.$
By (2.5), it suffices to show that for any $A_{n}\in\Sigma_{\mathcal{Y}}$, $A_{n-1}\in\Sigma_{{\mathcal{Y}}^{n-1}}$, $B\in\Sigma_{\Theta}$ we have
$\int_{B}{\mathbf{p}}_{x_{n}}(A_{n}|\theta){\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}(A_{k-1}|\theta)d\mu_{\Theta}(\theta)$
(3.9) $=\int_{A_{n-1}\times A_{n}}{\mathbf{q}}_{x_{n}}\big(B|y_{n}\|{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}(y_{1},\ldots,y_{n-1})\big)d(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}})_{*}\mu_{\Theta}(y_{1},\ldots,y_{n}).$
By Lemma 3.2, the Markov kernel $\overline{\underline{{\mathbf{p}}_{x_{n}}}\circ\underline{{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}}}:{\mathcal{Y}}^{n-1}\to{\mathcal{P}}({\mathcal{Y}})$ is a regular conditional probability measure for the joint distribution $(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}})_{*}\mu_{\Theta}\in{\mathcal{P}}({\mathcal{Y}}^{n})$ with respect to the projection $\Pi_{{\mathcal{Y}}^{n-1}}:{\mathcal{Y}}^{n}\to{\mathcal{Y}}^{n-1}$. Taking into account Composition Formula (2.1), we have
(3.10) $\overline{\underline{{\mathbf{p}}_{x_{n}}}\circ\underline{{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}}}=(\underline{{\mathbf{p}}_{x_{n}}})_{*}\circ{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}.$
Applying the disintegration formula, and taking into account (3.10), we verify (3.9) as follows
$\int_{B}{\mathbf{p}}_{x_{n}}(A_{n}|\theta){\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}(A_{n-1}|\theta)d\mu_{\Theta}(\theta)$
$\stackrel{{\scriptstyle?}}{{=}}\int_{A_{n-1}}\int_{A_{n}}{\mathbf{q}}_{x_{n}}\big(B|y_{n}\|{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}(y_{1},\ldots,y_{n-1})\big)d(\underline{{\mathbf{p}}_{x_{n}}})_{*}$
$\big({\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}(y_{1},\ldots,y_{n-1})\big)(y_{n})d(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}})_{*}\mu_{\Theta}(y_{1},\ldots,y_{n-1})$
(3.11) $=\int_{A_{n-1}}\int_{B}{\mathbf{p}}_{x_{n}}(A_{n}|\theta)d{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}(\theta|y_{1},\ldots,y_{n-1})d(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}})_{*}\mu_{\Theta}(y_{1},\ldots,y_{n-1}).$
The last equality holds since
${\mathbf{q}}_{x_{n}}\big(\cdot|{\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}(y_{1},\ldots,y_{n})\big):{\mathcal{Y}}\to{\mathcal{P}}(\Theta)$
is a Bayesian inversion of ${\mathbf{p}}_{x_{n}}:\Theta\to{\mathcal{Y}}$ relative to ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n-1})}(y_{1},\ldots,y_{n})\in{\mathcal{P}}(\Theta)$.
Since ${\mathbf{p}}_{x_{n}}(A_{n}|\cdot)\in{\mathcal{F}}_{b}(\Theta)$, fixing $A_{n-1}$ and $B$, we extend the LHS and RHS of (3.11) as linear functions on ${\mathcal{F}}_{b}(\Theta)$. Since ${\mathcal{F}}_{s}(\Theta)$ is dense in ${\mathcal{F}}_{b}(\Theta)$ in the sup-norm, to prove (3.9), it suffices to show that for any $B^{\prime}\in\Sigma_{\Theta}$ we have
$\int_{B\cap B^{\prime}}{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}(A_{n-1}|\theta)d\mu_{\Theta}(\theta)$
(3.12) $=\int_{A_{n-1}}\int_{B\cap B^{\prime}}d{\mathbf{q}}_{\Pi_{{\mathcal{X}}(S_{n-1})}}(\theta|y_{1},\ldots,y_{n-1})d(\underline{{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}})_{*}\mu_{\Theta}(y_{1},\ldots,y_{n-1}).$
Equation (3.12) holds since ${\mathbf{q}}_{\Pi_{{\mathcal{X}}(S_{n-1})}}:{\mathcal{Y}}^{n-1}\to{\mathcal{P}}(\Theta)$ is a Bayesian inversion of ${\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}:\Theta\to{\mathcal{Y}}^{n-1}$ relative to $\mu_{\Theta}$.
For a finite set $\mathcal{X}$} and a measurable space $\mathcal{Y}$}, we propose another method for computing Bayesian inversions on universal Bayesian models $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\\ \operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ using projective limits.
We denote by $\pi({\mathcal{Y}})$ the set of all finite partitions of $\mathcal{Y}$} into measurable subsets.
Denote by $|(A)|$ the size of a finite partition $(A)=\dot{\cup}_{i=1}^{k}A_{i}\in\pi({\mathcal{Y}})$, and by ${\Omega}_{(A)}:=\{A_{1},\ldots,A_{k}\}$ the finite set of size $|(A)|$ associated with $(A)$. Each partition $(A)=\dot{\cup}_{i=1}^{k}A_{i}\in\pi({\mathcal{Y}})$ is defined uniquely by a surjective measurable map
$\pi_{(A)}:{\mathcal{Y}}\to{\Omega}_{(A)}$
that maps $A_{i}\ni y$ to $A_{i}$. If a partition $(A)=(A_{1},\ldots,A_{k})$ of size $k$ of $\mathcal{Y}$} is a refinement of a partition $(A^{\prime})=(A^{\prime}_{1},\ldots,A_{l}^{\prime})$ of size $l\leq k$ we write $(A^{\prime})\leq(A)$. Then there exists a map
$\pi^{(A)}_{(A^{\prime})}:{\Omega}_{(A)}\to{\Omega}_{(A^{\prime})}$
such that
(3.13) $\pi_{(A^{\prime})}=\pi^{(A)}_{(A^{\prime})}\circ\pi_{(A)}.$
Thus $(\pi({\mathcal{Y}}),\leq)$ is a directed set of finite (measurable) partitions of $\mathcal{Y}$}.
Theorem 3.4.
Let $\mathcal{Y}$} be a measurable space, and ${\mathcal{X}}:=\{x_{1},\ldots,x_{n}\}$ a finite set and $X_{k}\in{\mathcal{X}}^{k}$. Let $\mu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$. Assume that for any $(A)\in\pi({\mathcal{Y}})$ there exists a Bayesian inversion ${\mathbf{q}}^{(A)}_{X_{k}}:{\Omega}_{(A)}^{k}\to{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{(A)})^{\mathcal{X}})$ of $\mathfrak{m}^{k}\circ E_{X_{k}}:{\mathcal{P}}({\Omega}_{(A)})^{\mathcal{X}}\to{\mathcal{P}}({\Omega}_{(A)}^{k})$ relative to $P_{*}(\pi_{(A)})_{*}^{\mathcal{X}}(\mu)\in{\mathcal{P}}({\mathcal{P}}({\Omega}_{(A)})^{\mathcal{X}})$ such that for any $(A)\leq(B)\in\pi({\mathcal{Y}})$ the following diagram is commutative:
${{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{(B)})^{\mathcal{X}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(\pi^{(B)}_{(A)})^{\mathcal{X}}_{*}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(B)}^{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}^{(B)}_{X_{k}}}$$\scriptstyle{(\pi^{(B)}_{(A)})^{k}}$${{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{(A)})^{\mathcal{X}}\big)}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}^{k}.}$$\scriptstyle{{\mathbf{q}}^{(A)}_{X_{k}}}$
Assume that there exists a map ${\mathbf{q}}_{X_{k}}:{\mathcal{Y}}^{k}\to{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ such that for any $(A)\in\pi({\mathcal{Y}})$ the following diagram is commutative for any $(A)\in\pi({\mathcal{Y}})$.
(3.14) ${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(\pi_{(A)})^{\mathcal{X}}_{*}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\mathcal{Y}}^{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}_{X_{k}}}$$\scriptstyle{(\pi_{(A)})^{k}}$${{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{(A)})^{\mathcal{X}}\big)}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}^{k}.}$$\scriptstyle{{\mathbf{q}}^{(A)}_{X_{k}}}$
Then ${\mathbf{q}}_{X_{k}}$ is a Bayesian inversion of $\mathfrak{m}^{k}\circ E_{X_{k}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}}^{k})$ relative to $\mu$.
Proof.
To prove Theorem 3.4, it suffices to show that ${\mathbf{q}}_{X_{k}}$ is a measurable map and for any ${\mathcal{A}}={\mathcal{A}}_{1}\times\ldots\times{\mathcal{A}}_{n}\in\Sigma_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}}$, ${\mathcal{A}}_{i}\in\Sigma_{{\mathcal{P}}({\mathcal{Y}})}$, $B=B_{1}\times\ldots\times B_{k}\in\Sigma_{{\mathcal{Y}}^{k}}$, $B_{j}\in\Sigma_{\mathcal{Y}}$, we have
(3.15) $(\Gamma_{\underline{{\mathbf{q}}_{X_{k}}}})_{*}(\underline{\mathfrak{m}^{k}\circ E_{X_{k}}})_{*}\mu(B\times{\mathcal{A}})=(\Gamma_{\underline{\mathfrak{m}^{k}\circ E_{X_{k}}}})_{*}\mu({\mathcal{A}}\times B).$
Recall that $\Sigma_{{\mathcal{P}}({\mathcal{Y}})}$ is generated by subsets $e^{-1}_{A}(C)$ where $A\in\Sigma_{\mathcal{Y}}$, $C\in{\mathcal{B}}({\mathbb{R}})$, and $e_{A}:{\mathcal{P}}({\mathcal{Y}})\to{\mathbb{R}}$ is defined by $\mu\mapsto\mu(A)$.
Lemma 3.5.
Let ${\mathcal{A}}_{i}=e^{-1}_{A_{i}}(C_{i})\in\Sigma_{{\mathcal{P}}({\mathcal{Y}})}$ for $i\in\overline{1,n}$. Then there exists a finite partition $(\tilde{A})\in\pi({\mathcal{Y}})$ and a subset ${\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{n}\in\Sigma_{{\mathcal{P}}\big({\Omega}_{(\tilde{A})}\big)}$ such that
(3.16) $\big((\pi_{(\tilde{A})})_{*}\big)^{-1}({\mathcal{S}}_{i})={\mathcal{A}}_{i}\text{ for all }i\in\overline{1,n}.$
Hence, $\Sigma_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}}$ is generated by subsets $\Pi_{j=1}^{n}\big(\pi_{(A)})_{*}^{-1}(ev_{S_{i}^{j}}^{-1}(C_{i}^{j}))$ where $C^{j}_{i}\in{\mathcal{B}}({\mathbb{R}})$ and $S_{i}^{j}\subset{\Omega}_{(A)}$, $(A)\in\pi({\mathcal{Y}})$.
Proof of Lemma 3.5.
Let $(\tilde{A})\in\pi({\mathcal{Y}})$ be a finite partition such that for any $i\in\overline{1,n}$ there exists a subset $S_{i}\in{\Omega}_{|(\tilde{A})|}$ such that
(3.17) $\pi_{(\tilde{A})}^{-1}(S_{i})=A_{i}.$
Then we have the following commutative diagram for any $i\in\overline{1,n}$:
[…]
It follows that ${\mathcal{A}}_{i}=e_{A_{i}}^{-1}(C_{i})={\mathcal{S}}_{i}:=\Big(\big(\pi_{(\tilde{A})}\big)_{*}\Big)^{-1}(e^{-1}_{S_{i}}(C_{i}))$. This proves (3.16). The last assertion of Lemma 3.5 follows immediately. ∎
Completion of the proof of Theorem 3.4. The measurability of ${\mathbf{q}}_{X_{k}}$ follows from Lemma 3.5, taking into account the commutative diagram 3.14. To complete the proof of Theorem 3.4, we need to prove (3.15). By Lemma 3.5, the RHS of (3.15) equals $(\Gamma_{\underline{\mathfrak{m}^{k}\circ E_{X_{k}}}}P_{*}(\pi_{(\tilde{A})})^{\mathcal{X}}_{*}\mu({\mathcal{S}}_{1}\times\ldots\times{\mathcal{S}}_{n}\times B)$ and the LHS of (3.15) equals $(\Gamma_{\underline{{\mathbf{q}}^{(\tilde{A})}_{X_{k}}}})_{*}(\underline{\mathfrak{m}^{k}\circ E_{X_{k}}})_{*}P_{*}(\pi_{(\tilde{A})})^{\mathcal{X}}_{*}\mu$. Since ${\mathbf{q}}^{(\tilde{A})}_{X_{k}}$ is a Bayesian inversion of $\mathfrak{m}^{k}\circ E_{X_{k}}$, (3.15) holds. ∎
Example 3.6 (Posterior distributions of Dirichlet processes).
For a measurable space $\mathcal{Y}$} denote by ${\mathcal{M}}^{*}({\mathcal{Y}})$ the measurable space of all non-zero finite measures on $\mathcal{Y}$} whose $\sigma$-algebra is defined in the same way as the $\sigma$-algebra $\Sigma_{w}$ on ${\mathcal{P}}({\mathcal{Y}})$, see [JLT21, §2.1]. By [JLT21, Theorem 4], there exists a measurable map ${\mathcal{D}}:{\mathcal{M}}^{*}({\mathcal{Y}})\to{\mathcal{P}}^{2}({\mathcal{Y}})$ such that ${\mathcal{D}}(\alpha)$ is the Dirichlet measure on ${\mathcal{P}}({\mathcal{Y}})$ with parameter $\alpha$ and for any $\alpha\in{\mathcal{M}}^{*}({\mathcal{Y}})$ the following diagram is commutative
[…]
Here $Dir_{(A)}(\beta)$ is the Dirichlet distribution with parameter $\beta\in{\mathcal{M}}^{*}({\Omega}_{(A)})$, and $M_{*}(\pi_{(A)}):{\mathcal{M}}^{*}({\mathcal{Y}})\to{\mathcal{M}}({\Omega}_{{}_{|(A)|}})$ is defined by the same formula (2.2). Let us consider a Bayesian statistical model $\big({\mathcal{P}}({\mathcal{Y}}),{\mathcal{D}}(\alpha),\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})},{\mathcal{Y}}\big)$ associated with the case #(X)=1\#(X)=1 and $k=1$ in Theorem 3.4. Lemma 3.5 and the above commutative diagram imply that the condition of Theorem 3.4 holds for $\big({\mathcal{P}}({\mathcal{Y}}),{\mathcal{D}}(\alpha),\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})},{\mathcal{P}}({\mathcal{Y}})\big)$. It is known that
${\mathbf{q}}_{(A)}(\cdot\|Dir(M_{*}(\pi_{(A)})\alpha):{\Omega}_{(A)}\to{\mathcal{P}}^{2}({\Omega}_{(A)}),x\mapsto Dir(M_{*}(\pi_{(A)})\alpha+\delta_{x})$
is a Bayesian inversion of the Markov kernel $\operatorname{Id}_{{\mathcal{P}}({\Omega}_{(A)})}$ relative to ${\mathcal{D}}\big(M_{*}(\pi_{(A)}\alpha)\big)$.
For $(A)\leq(B)\in(\pi({\mathcal{Y}}),\leq)$, we verify immediately that the following diagram
${{\mathcal{P}}^{2}({\Omega}_{(B)})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P^{2}_{*}\pi^{(B)}_{(A)}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(B)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}_{(B)}(\cdot\|{\mathcal{D}}(M_{*}(\pi_{(B)}\alpha)))}$$\scriptstyle{\pi^{(B)}_{(A)}}$${{\mathcal{P}}^{2}({\Omega}_{(A)})}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}.}$$\scriptstyle{{\mathbf{q}}_{(A)}(\cdot\|{\mathcal{D}}(M_{*}(\pi_{(A)}\alpha)))}$
is commutative. Hence, by Theorem 3.4, the map
${\mathbf{q}}:{\mathcal{Y}}\to{\mathcal{P}}^{2}({\mathcal{Y}}),y\mapsto{\mathcal{D}}(\alpha+\delta_{y})$
is a Bayesian inversion of $\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})}$ relative to ${\mathcal{D}}(\alpha)$. Taking into account Theorem 3.1, the map
${\mathbf{q}}^{n}:{\mathcal{Y}}^{n}\to{\mathcal{P}}^{2}({\mathcal{Y}}),(y_{1},\ldots,y_{n})\mapsto{\mathcal{D}}(\alpha+\sum_{i=1}^{n}\delta_{y_{i}})$
is a Bayesian inversion of the Markov kernel
$\mathrm{\operatorname{Id}}^{n}_{{\mathcal{P}}({\mathcal{Y}})}:{\mathcal{P}}({\mathcal{Y}})\to{\mathcal{P}}({\mathcal{Y}}^{n}),\mu\mapsto\otimes^{n}\mu$
relative to ${\mathcal{D}}(\alpha)$.
4. Posterior predictive distributions
In this Section, using Theorem 3.1, we shall prove Theorem 4.1 and 4.3 on recursive computing posterior predictive distributions in Bayesian supervised learning.
Theorem 4.1 (Posterior predictive distribution).
Let $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ be a Bayesian model for supervised learning, $X_{n}=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$, and $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$. For $Y_{n}=(y_{1},\ldots,y_{n})\in{\mathcal{Y}}^{n}$ let $S_{n}(X_{n},Y_{n}):=\big((x_{1},y_{1}),\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$.
1) The posterior predictive distribution ${\mathcal{P}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ defined by Equation (2.10) is a regular conditional probability measure for the joint distribution
$\mu^{0}_{T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta}}:=\underline{\mathfrak{m}^{2}({\mathbf{p}}_{T_{m}},{\mathbf{p}}_{X_{n}})}_{*}\mu_{\Theta}\in{\mathcal{P}}({\mathcal{Y}}^{m}\times{\mathcal{Y}}^{n})$
with respect to the projection $\Pi_{{\mathcal{Y}}^{n}}:{\mathcal{Y}}^{m}\times{\mathcal{Y}}^{n}\to{\mathcal{Y}}^{n}$. Hence, if ${\mathbf{q}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ is a regular conditional probability measure for $\mu^{0}_{T_{m},S_{n}(X_{m}Y_{n}),\mu_{\Theta}}$ with respect to the projection $\Pi_{{\mathcal{Y}}^{n}}$ then ${\mathbf{q}}^{n}_{m}\big(Y_{n}\big)$ coincides with ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ in the weak sense up to a $(\underline{{\mathbf{p}}_{X_{n}}})_{*}\mu_{\Theta}$-zero set.
2) The posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ can be computed recursively as follows.
  1. (1)
    Step 1: Let ${\mathbf{q}}^{1}_{m+n-1}:{\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}^{m+n-1})$ be a regular conditional probability measure for the joint distribution $\mu^{0}_{T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}\in{\mathcal{P}}({\mathcal{Y}}^{m+n})$ with respect to the projection ${\mathcal{Y}}^{m+n}\to{\mathcal{Y}}$. Then we set
    (4.1) $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{1}:={\mathbf{q}}^{1}_{m+n-1}(y_{n})\in{\mathcal{P}}({\mathcal{Y}}^{m+n-1}).$
  2. (2)
    Step $k+1$ for $1\leq k\leq n-1$. Let ${\mathbf{q}}^{k+1}:{\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}^{m+n-k-1})$ be a regular conditional probability measure for the joint distribution $\mu^{k}_{T_{m},S_{n},\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m+n-k})$ with respect to the projection ${\mathcal{Y}}^{m+n-k}\to{\mathcal{Y}}$. Then we set
    (4.2) $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{k+1}:={\mathbf{q}}^{k}(y_{n-k})\in{\mathcal{P}}({\mathcal{Y}}^{m+n-k-1}).$
Then $\mu_{T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta}}^{n}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ is the posterior predictive distribution of ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}$.
Proof.
1) The first assertion of Theorem 4.1 is a direct consequence of Proposition 3.3.
2) For $1\leq k\leq n$ we let $S_{k}:=\big((x_{1},y_{1}),\ldots,(x_{k},y_{k})\big)$. To prove the second assertion of Theorem 4.1, we consider the following diagram
[…]
where
${\mathbf{p}}_{m+n-1}:=\mathfrak{m}^{2}({\mathbf{p}}_{T_{m}},{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-1})}),\,{\mathbf{p}}_{m+n}:=\mathfrak{m}^{2}({\mathbf{p}}_{T_{m}},{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}).$
By Proposition 3.3 we have
$\mu_{(T_{m},S_{n},\mu_{\Theta})}^{1}={\mathcal{P}}_{T_{m}|(x_{n},y_{n}),\mu_{\Theta}}.$
Next, we consider the following diagram
${\ignorespaces\ignorespaces\ignorespaces\ignorespaces(\Theta,\mu_{\Theta|(x_{n},y_{n})})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\underline{{\mathbf{p}}_{m+n-2}}}$$\scriptstyle{\underline{{\mathbf{p}}_{m+n-1}}}$$\scriptstyle{\underline{{\mathbf{p}}_{x_{n-1}}}}$${{\mathcal{Y}}^{m+n-2}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\mathcal{Y}}^{m+n-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi_{{\mathcal{Y}}^{m+n-2}}}$$\scriptstyle{\Pi_{\mathcal{Y}}}$${\mathcal{Y}}$
where
${\mathbf{p}}_{m+n-2}:=\mathfrak{m}^{2}({\mathbf{p}}_{T_{m}},{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n-2})}).$
Applying Proposition 3.3, taking into account Theorem 3.1, we obtain
$\mu_{(T_{m},S_{n},\mu_{\Theta})}^{2}={\mathcal{P}}_{T_{m}|(x_{n},y_{n}),(x_{n-1},y_{n-1}),\mu_{\Theta}}.$
Repeating this procedure, we obtain
$\mu^{n}_{(T_{m},S_{n},\mu_{\Theta})}={\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}.$
For $X_{m}:=(x_{1},\ldots,x_{m})\in{\mathcal{X}}^{m}$ we denote by $X_{m}]$ the smallest subset of $\mathcal{X}$} that contains each of $x_{i}$.
Corollary 4.2.
Let $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ be a universal Bayesian supervised models. Let $T_{m}\in{\mathcal{X}}^{m},X_{n}\in{\mathcal{X}}^{n}$ and $A:=[T_{m}]\cup[X_{n}]$. Let $R_{A}^{\mathcal{X}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{A},h\mapsto h_{|A},$ denote the restriction map. Denote by $\operatorname{Id}_{X_{n}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}}^{n})$ the composition $\mathfrak{m}^{n}\circ E_{X_{n}}$. Then for $(\underline{\operatorname{Id}_{X_{n}}})_{*}\mu$-a.s. $Y_{n}\in{\mathcal{Y}}^{n}$ we have
(4.3) ${\mathcal{P}}_{T_{m}|S_{n}(X_{,}Y_{n}),\mu}={\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),(R_{A}^{\mathcal{X}})_{*}\mu}$
in the weak sense, where the RHS of (4.3) is the posterior predictive distribution of the restricted Bayesian supervised learning model $({\mathcal{P}}({\mathcal{Y}})^{A},(R_{A}^{\mathcal{X}})_{*}\mu_{\Theta},\\ \operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{A}},{\mathcal{P}}({\mathcal{Y}})^{A})$.
Proof.
We consider the following diagram
[…]
where $\operatorname{Id}_{T_{m}}:=\mathfrak{m}^{m}\circ E_{T_{m}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ and $\operatorname{Id}^{A}_{T_{m}}:=\mathfrak{m}^{m}\circ E_{T_{m}}:{\mathcal{P}}({\mathcal{Y}})^{A}\to{\mathcal{P}}({\mathcal{Y}}^{m})$. Similarly, we define $\operatorname{Id}^{A}_{X_{n}}$. By Theorem 4.1, ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},\cdot),\mu}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ is a regular conditional probability measure for the joint distribution $(\underline{\operatorname{Id}_{T_{m},X_{n}}})_{*}\mu\in{\mathcal{P}}({\mathcal{Y}}^{m+n})$, and ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},\cdot),(R_{A}^{\mathcal{X}})_{*}\mu}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ is a regular conditional probability measure for the joint distribution $(\underline{\operatorname{Id}^{A}_{T_{m},X_{n}}})_{*}\\ (R_{A}^{\mathcal{X}})_{*}\mu\in{\mathcal{P}}({\mathcal{Y}}^{m+n})$. To conclude Corollary 4.2, we note that
$(\underline{\operatorname{Id}_{T_{m},X_{n}}})_{*}\mu=(\underline{\operatorname{Id}^{A}_{T_{m},X_{n}}})_{*}(R_{A}^{\mathcal{X}})_{*}\mu$
since $T_{m}]\subset A$ and $X_{n}]\subset A$, hence $\operatorname{Id}_{T_{m},X_{n}}=\operatorname{Id}^{A}_{T_{m},X_{n}}\circ R_{A}^{\mathcal{X}}$. Therefore, both the Bayesian statistical supervised model $({\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}})$ and the Bayesian statistical supervised model $({\mathcal{P}}({\mathcal{Y}}^{A},(R^{\mathcal{X}}_{A})_{*}\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{A}})$ induce the same joint distribution on ${\mathcal{P}}({\mathcal{Y}}^{m+n})$. Taking into account Theorem 4.1, we obtain Corollary 4.2. ∎
Let us now consider Bayesian regression learning, which is a particular case of Bayesian supervised learning [Le2025, Definition 3.10]. Let $\mathcal{X}$} be an input space and $V={\mathbb{R}}^{k}$. We consider a corrupted measurement
(4.5) $y=f(x)+{\varepsilon}(x)\in V,\qquad f\in(V^{\mathcal{X}},\mu),\qquad{\varepsilon}(x)\in(V,\nu_{\varepsilon}(x))$
where $\nu_{\varepsilon}(x)\in{\mathcal{P}}(V)$ for all $x\in{\mathcal{X}}$. We regard $V^{\mathcal{X}}$ as a universal parameter space, and the quadruple $(V^{\mathcal{X}},\mu,{\mathbf{p}}^{\varepsilon},{\mathcal{P}}(V)^{\mathcal{X}})$ with
${\mathbf{p}}^{\varepsilon}(f):=\delta_{f}*\nu_{\varepsilon},$
(4.6) $\delta_{f}*\nu_{\varepsilon}(x):=\delta_{f(x)}*\nu_{\varepsilon}(x),$
as a universal Bayesian supervised learning model for learning the corrupted measurements (4.5).
In the general case, we consider a quadruple $(\Theta,\mu_{\Theta},h,V^{\mathcal{X}})$ where $(\Theta,\mu_{\Theta})$ is a parameter space with a prior probability measure $\mu_{\Theta}$, and $h\in{\bf Meas}(\Theta,V^{\mathcal{X}})$. For $X_{n}=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$, the Markov kernel $\mathfrak{m}^{n}\circ E_{X_{n}}\circ{\mathbf{p}}^{\varepsilon}\circ h:\Theta\to{\mathcal{P}}(V^{n})$ describes the sampling distribution of the joint distribution of $(y_{1},\ldots,y_{n})$ where $y_{i}=f(x_{i})+{\varepsilon}(x_{i})$.
Let
${\mathbf{p}}^{0}:V^{\mathcal{X}}\to{\mathcal{P}}(V^{\mathcal{X}}),f\mapsto\delta_{f},$
be the Markov kernel describing the sampling distribution of uncorrupted measurement. By Proposition 3.3, for $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$, the predictive distribution of the tuple $(f(t_{1}),\ldots,f(t_{m}))$ after seeing $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$ can be chosen as the value ${\mathbf{q}}^{n}_{m}(\Pi_{\mathcal{Y}}(S_{n}))\in{\mathcal{P}}(V^{m})$ where ${\mathbf{q}}^{n}_{m}:V^{m}\to V^{n}$ is a regular conditional probability measure for the joint distribution $(\underline{\mathfrak{m}^{2}(\mathfrak{m}^{m}\circ E_{T_{m}}\circ{\mathbf{p}}^{0}\circ h,\mathfrak{m}^{n}\circ E_{\Pi_{\mathcal{X}}(S_{n})}\circ{\mathbf{p}}^{\varepsilon}\circ h)})_{*}\mu_{\Theta}\in{\mathcal{P}}(V^{m}\times V^{n})$.
We shall abbreviate $\mathfrak{m}^{m}\circ E_{T_{m}}\circ{\mathbf{p}}^{0}\circ h$ as $h^{0}_{T_{m}}$, and $\mathfrak{m}^{m}\circ E_{T_{m}}\circ{\mathbf{p}}^{\varepsilon}\circ h$ as $h^{\varepsilon}_{T_{m}}$. The following theorem for Bayesian regression learning is proved in the same way as Theorem 4.1, so we omit its proof.
Theorem 4.3 (Posterior predictive distribution with corrupted measurements).
Let $(\Theta,\mu_{\Theta},h,V^{\mathcal{X}})$ be a Bayesian model for regression learning, with $y_{i}$ being a corrupted measurement of $f(x_{i})$ for $i\in\overline{1,n}$, and $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$. For $X_{n}=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$ and $Y_{n}=(y_{1},\ldots,y_{n})\in V^{n}$ let $S_{n}(X_{n},Y_{n}):=\big((x_{1},y_{1}),\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$.
1) The posterior predictive distribution
${\mathcal{P}}^{n}_{m}:V^{n}\to{\mathcal{P}}(V_{m}),\,Y_{n}\mapsto{\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}},$
is a regular conditional probability measure for the joint distribution
$\mu^{0}_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}:=\underline{\mathfrak{m}^{2}(h^{0}_{T_{m}},h^{\varepsilon}_{\Pi_{\mathcal{X}}(S_{n})})}_{*}\mu_{\Theta}\in{\mathcal{P}}(V^{m}\times V^{n})$
with respect to the projection $\Pi_{V^{n}}:V^{m}\times V^{n}\to V^{n}$, Hence, if ${\mathbf{q}}^{n}_{m}:V^{n}\to{\mathcal{P}}(V^{m})$ is a regular conditional probability measure for the joint distribution $\mu^{0}_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}$ with respect to the projection $\Pi_{V^{n}}$ then ${\mathbf{q}}^{n}_{m}\big(\Pi_{V}(S_{n})\big)$ coincides with ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$ in the weak sense up to a $(\underline{h^{\varepsilon}_{X_{n}}})_{*}\mu_{\Theta}$-zero set.
2) The posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$ can be computed recursively as follows.
  1. (1)
    Step 1: Let ${\mathbf{q}}^{1}:V\to{\mathcal{P}}(V^{m+n-1})$ be a regular conditional probability measure for the joint distribution $\mu^{0}_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}\in{\mathcal{P}}(V^{m+n})$ with respect to the projection $V^{m+n}\to V$. Then we set
    (4.7) $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{1}:={\mathbf{q}}^{1}(y_{n})\in{\mathcal{P}}(V^{m+n-1}).$
  2. (2)
    Step $k+1$ for $1\leq k\leq n-1$. Let ${\mathbf{q}}^{k+1}:V\to{\mathcal{P}}(V^{m+n-k-1})$ be a regular conditional probability measure for the joint distribution $\mu^{k}_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}\in{\mathcal{P}}(V^{m+n-k})$ with respect to the projection $V^{m+n-k}\to V$. Then we set
    (4.8) $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{k+1}:={\mathbf{q}}^{k}(y_{n-k})\in{\mathcal{P}}(V^{m+n-k-1}).$
Then $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{n}\in{\mathcal{P}}(V^{m})$ is the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$.
Example 4.4 (Gaussian process regression).
We illustrate Theorem 4.3 with Gaussian process regression model $({\mathbb{R}}^{\mathcal{X}},{\mathcal{G}}{\mathcal{P}}(m,K),{\mathbf{p}}^{\varepsilon},{\mathcal{P}}({\mathbb{R}})^{\mathcal{X}})$, where ${\mathcal{G}}{\mathcal{P}}(m,k)$ is a Gaussian measure on $R^{\mathcal{X}}$ defined by its mean function $m\in{\mathbb{R}}^{\mathcal{X}}$ and its covariance function $K:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$, which is a positive definite kernel. Let $\mu$ be a Gaussian measure on the function space $R^{\mathcal{X}}$ where $\mathcal{X}$} is an input space and $V={\mathbb{R}}$. One sees immediately that the recipe for computing the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},{\mathcal{G}}{\mathcal{P}}(m,K)}$ in Theorem 4.3(1) coincides with the classical formula for posterior predictive distributions in Gaussian process regression described in [RW2006]. Furthermore, the recursive formula in Theorem 4.3(2) is much simpler and faster than the classical formula since it does not require computing the (pseudo) inverse of a square matrix of size $(n\times n)$ associated with the kernel $K^{\varepsilon}_{n}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ which is the variance of the Gaussian measure that governs the distribution of $y_{i}=f(x_{i})+{\varepsilon}(x_{i})\in{\mathbb{R}}$, $i\in\overline{1,n}$, and a multiplication of matrices of size $(m\times n)$ with this pseudo inverse and with a matrix of size $(n\times m)$ [Stein1999, Appendix A], [Rao2002, Chapter 8]. Instead, we have to compute $n$-round of multiplications of matrices of size $m+n-i\times 1$ with a matrix of size $1\times m+n-i$ for $i\in\overline{1,n}$. This sequential update procedure is known to be equivalent to the celebrated Kalman filter update equations, see, e.g., [SS2023, §6.3].
5. Probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$
In this Section, we assume that $\mathcal{Y}$} is a Polish space unless otherwise stated. Then ${\mathcal{P}}({\mathcal{Y}})^{k}$ and ${\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{k})$ are Polish spaces for any $k\in{\mathbb{N}}^{+}$. Using Orbanz’s description of the space ${\mathcal{P}}^{2}({\mathcal{Y}})$ for a Polish space $\mathcal{Y}$} [Orbanz2011, Theorem 1.1], we shall describe the space ${\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ using a projective system (Theorem 5.5).
For a set $\mathcal{X}$} we denote by $\mathrm{P_{fin}}({\mathcal{X}})$ the directed set of finite subsets of $\mathcal{X}$}. Our projective system is a product of two projective systems. The first projective system is associated to the restriction maps
$R_{X_{m}}^{X_{n}}:{\mathcal{P}}({\mathcal{Y}})^{X_{n}}\to{\mathcal{P}}({\mathcal{Y}})^{X_{m}}\text{ if }X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}}).$
Recall that $R^{\mathcal{X}}_{X_{m}}$ denotes the restriction map ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{{\mathcal{X}}_{m}}$.
Lemma 5.1.
Let $\mathcal{Y}$} be a Polish space. Then for any set $\mathcal{X}$} and $\mu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ we have
(5.1) $\mu=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}(R^{\mathcal{X}}_{X_{m}})_{*}\mu.$
Conversely, for any projective system of probability spaces $\{({\mathcal{P}}({\mathcal{Y}})^{X_{m}},\mu_{X_{m}}),R_{X_{m}}^{X_{n}}:X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})\}$ there exists a unique probability measure $\mu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that for all $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ we have
$\mu_{X_{m}}=(R^{\mathcal{X}}_{X_{m}})_{*}\mu.$
Proof.
The first assertion is straightforward. The second assertion is obtained by applying the Kolmogorov extension theorem. ∎
Next we shall study another projective system associated with partitions of a Polish space $\mathcal{Y}$}.
Let ${\mathcal{A}}_{\mathcal{Y}}$ be the countable algebra generated by open balls with rational radius centered at a countable dense set in $\mathcal{Y}$}. Then ${\mathcal{A}}_{\mathcal{Y}}$ generates the Borel $\sigma$-algebra of $\mathcal{Y}$}. Set
${\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}}):=\{(A):=(A_{1},\ldots,A_{n}):A_{i}\in{\mathcal{A}}_{\mathcal{Y}},\,\dot{\cup}A_{i}={\mathcal{Y}}\}.$
Note that $\big({\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}}),\leq\big)$ is a directed subset of $\big(\pi({\mathcal{Y}}),\leq\big)$.
Remark 5.2.
Using (3.13) one observes that the collection
$\{{\mathcal{P}}({\Omega}_{(A)}),(\pi^{(A)}_{(B)})_{*}:{\mathcal{P}}({\Omega}_{(A)})\to{\mathcal{P}}({\Omega}_{(B)}),\,(B)\leq(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})\}$
forms a projective system of topological spaces. Since ${\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ is countable, by Bochner theorem [Bochner1955], [Bourbaki2004, Theorem 2, chapter IX, §4.3], [Orbanz2011, Theorem 2.2] there exists the projective limit
${\mathcal{C}}_{{\mathcal{A}}_{\mathcal{Y}}}:=\lim_{\stackrel{{\scriptstyle\longleftarrow}}{{(A)\in{\mathcal{H}}(A_{\mathcal{Y}})}}}(\pi_{(A)})_{*}{\mathcal{P}}({\mathcal{Y}})$
in the category of topological spaces.
Let ${\mathcal{P}}_{{\mathcal{A}}_{\mathcal{Y}}}$ denote the set of all probability measures on ${\mathcal{A}}_{\mathcal{Y}}$, i.e.,
${\mathcal{P}}_{{\mathcal{A}}_{\mathcal{Y}}}:={\mathcal{P}}({\mathcal{Y}})_{|{\mathcal{A}}_{\mathcal{Y}}}.$
By the Caratheodory extension theorem, each $\mu\in{\mathcal{P}}({\mathcal{Y}})$ is defined uniquely by its restriction $\mu_{|{\mathcal{A}}_{\mathcal{Y}}}$. Clearly we have ${\mathcal{P}}_{{\mathcal{A}}_{\mathcal{Y}}}\subset{\mathcal{C}}_{{\mathcal{A}}_{\mathcal{Y}}}$.
In [Orbanz2011, Proposition 3.1] Orbanz proved the following.
Proposition 5.3.
(1) ${\mathcal{C}}_{{\mathcal{A}}_{\mathcal{Y}}}$ is the space of all probability charges on ${\mathcal{A}}_{\mathcal{Y}}$.
(2) ${\mathcal{P}}_{{\mathcal{A}}_{\mathcal{Y}}}$ is a measurable subset of $\big({\mathcal{C}}_{{\mathcal{A}}_{\mathcal{Y}}},{\mathcal{B}}({\mathcal{C}}_{{\mathcal{A}}_{\mathcal{Y}}})\big)$.
(3) The restriction map $R_{{\mathcal{A}}_{\mathcal{Y}}}:{\mathcal{P}}({\mathcal{Y}})\to\big({\mathcal{P}}_{{\mathcal{A}}_{\mathcal{Y}}},{\mathcal{B}}({\mathcal{C}}_{{\mathcal{A}}_{\mathcal{Y}}})\cap{\mathcal{P}}_{{\mathcal{A}}_{\mathcal{Y}}}\big)$ is an isomorphism of measurable spaces.
For a measurable space $\mathcal{Y}$} and $k\in{\mathbb{N}}^{+}$, denote by
$\widehat{e}v_{\mathcal{Y}}^{(k)}:{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{k}\big)\to{\mathcal{P}}({\mathcal{Y}}^{k})$
the marginalization map,
(5.2) $\widehat{ev}_{\mathcal{Y}}^{(k)}(\nu)(B_{1}\times\ldots\times B_{k}):=\int_{{\mathcal{P}}({\mathcal{Y}})^{k}}\mathfrak{m}^{k}\mu(B_{1}\times\ldots\times B_{k})\,d\nu(\mu)$
for $\nu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{k}\big)$ and $B_{i}\in\Sigma_{\mathcal{Y}}$, $i\in\overline{1,k}$. Similarly, for a finite set $X_{m}:=\{x_{1},\ldots,x_{m}\}$ we denote by
$\widehat{e}v_{\mathcal{Y}}^{X_{m}}:{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{m}}\big)\to{\mathcal{P}}({\mathcal{Y}}^{X_{m}})$
the marginalization map,
(5.3) $\widehat{ev}_{\mathcal{Y}}^{X_{m}}(\nu)(B_{1}\times\ldots\times B_{m}):=\int_{{\mathcal{P}}({\mathcal{Y}})^{X_{m}}}\mathfrak{m}^{m}\mu(B_{1}\times\ldots\times B_{m})\,d\nu(\mu)$
for $\nu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{m}}\big)$ and $B_{i}\in\Sigma_{{\mathcal{Y}}^{\{x_{i}\}}}$, $i\in\overline{1,m}$.
Lemma 5.4.
1) Recall that $\mathfrak{m}^{k}:{\mathcal{P}}({\mathcal{Y}})^{k}\to{\mathcal{P}}({\mathcal{Y}}^{k})$ is the multiplication map. Then we have
(5.4) $\widehat{ev}^{(k)}_{\mathcal{Y}}=P_{*}\underline{\mathfrak{m}^{k}}.$
2) The map $\widehat{ev}_{\mathcal{Y}}^{(k)}$ is measurable. If $\mathcal{Y}$} is a Polish space, then $\widehat{ev}_{\mathcal{Y}}^{(k)}$ is a $(\tau_{w},\tau_{w})$-continuous mapping.
3) Assume that $\mathcal{Y}$} is a Polish space. Given a set $\mathcal{X}$} and $\nu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ the following formula for the marginalization $\widehat{ev}_{\mathcal{Y}}^{\mathcal{X}}(\nu)$ of $\nu$,
(5.5) $\widehat{ev}_{\mathcal{Y}}^{\mathcal{X}}(\nu):=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}\widehat{ev}_{\mathcal{Y}}^{X_{m}}\big((R^{\mathcal{X}}_{X_{m}})_{*}\nu\big)\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}}),$
is well-defined.
Proof.
(1) Equation (5.4) follows directly from (5.2) and Equation (2.2).
(2) The second assertion follows from the first one, taking into account the measurability of $P_{*}\underline{\mathfrak{m}^{k}}$ (and its $(\tau_{w},\tau_{w})$-continuity if $\mathcal{Y}$} is Polish) by [Giry82, Theorem1].
3) For $X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$ let $r^{\mathcal{X}}_{X_{m}}:{\mathcal{Y}}^{\mathcal{X}}\to{\mathcal{Y}}^{X_{m}}$ and $r^{X_{n}}_{X_{m}}:{\mathcal{Y}}^{X_{n}}\to{\mathcal{Y}}^{X_{m}}$ denote the restriction maps.
By definition of the multiplication map $\mathfrak{m}^{X_{n}}:{\mathcal{P}}({\mathcal{Y}})^{X_{n}}\to{\mathcal{P}}({\mathcal{Y}}^{X_{n}})$, we verify immediately that the following diagram
(5.6) ${{\mathcal{P}}({\mathcal{Y}})^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R^{X_{n}}_{X_{m}}}$$\scriptstyle{\underline{\mathfrak{m}^{X_{n}}}}$${{\mathcal{P}}({\mathcal{Y}})^{X_{m}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\underline{\mathfrak{m}^{X_{m}}}}$${{\mathcal{Y}}^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r^{X_{n}}_{X_{m}}}$${{\mathcal{Y}}^{X_{m}}}$
is commutative. Consequently, the following diagram for any $X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$:
(5.7) ${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{n}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(R^{X_{n}}_{X_{m}})}$$\scriptstyle{P_{*}\underline{\mathfrak{m}^{X_{n}}}}$${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{m}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}\underline{\mathfrak{m}^{X_{m}}}}$${{\mathcal{P}}({\mathcal{Y}}^{X_{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(r^{X_{n}}_{X_{m}})_{*}}$${{\mathcal{P}}({\mathcal{Y}}^{X_{m}})}$
is commutative. Combining the commutativity of the diagram (5.7), Equation (5.4), and the second assertion of Lemma 5.4, we complete the proof of Lemma 5.4, taking into account the Kolmogorov extension theorem. ∎
Let $\mathcal{Y}$} be a Polish space and $\mathcal{X}$} a set. For $(B)\leq(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$, and for $X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$ we denote by
$R^{X_{n}}_{(A),X_{m}}:{\Omega}_{(A)}^{X_{n}}\to{\Omega}_{(A)}^{X_{m}}$
the restriction map, and by
$\pi^{(A),X_{m}}_{(B)}:{\Omega}_{(A)}^{X_{m}}\to{\Omega}_{(B)}^{X_{m}}$
the natural projection map.
Theorem 5.5.
Assume that $\mathcal{Y}$} is a Polish space and ${\mathcal{A}}_{\mathcal{Y}}$ is the countable algebra generating ${\mathcal{B}}({\mathcal{Y}})$ defined above. Then for any $\nu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ we have
(5.8) $\nu=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}\lim_{\stackrel{{\scriptstyle\longleftarrow}}{{(A)\in{\mathcal{H}}(A_{\mathcal{Y}})}}}P^{2}_{*}\pi_{(A)}^{X_{m}}(R^{\mathcal{X}}_{X_{m}})_{*}\nu$
and
(5.9) $\widehat{ev}_{{\mathcal{Y}}}^{\mathcal{X}}(\nu)=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}\lim_{\stackrel{{\scriptstyle\longleftarrow}}{{(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})}}}P_{*}\pi_{(A)}^{X_{m}}(r^{\mathcal{X}}_{X_{m}})_{*}\widehat{ev}_{{\mathcal{Y}}}^{\mathcal{X}}(\nu).$
Conversely, given a projective system of finite sample spaces endowed with second order probability measures
$\Big\{\Big({\Omega}^{X_{m}}_{(A)},\nu_{(A)}^{X_{m}}\in{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{(A)})^{X_{m}}\big)\Big):X_{m}\in\mathrm{P_{fin}}({\mathcal{X}}),A\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})\Big\}$
and induced projection maps
$\Big\{(R_{(A),X_{m}}^{X_{n}})_{*}:{\mathcal{P}}({\Omega}_{(A)})^{X_{n}}\to{\mathcal{P}}({\Omega}_{(A)})^{X_{m}},$
$P_{*}(R_{(A),X_{m}}^{X_{n}})_{*}:{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{|A|})^{X_{n}}\big)\to{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{(A)})^{X_{m}}\big),$
$P_{*}^{2}\pi^{(A),X_{m}}_{(B)}:{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{(A)})^{X_{m}}\big)\to{\mathcal{P}}\big({\mathcal{P}}({\Omega}_{(B)})^{X_{m}}\big)|\colon$
$X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}}),\,(B)\leq(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})\Big\}$
there exists $\nu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that for any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ and $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ we have
(5.10) $\nu_{(A)}^{X_{m}}=P^{2}_{*}\pi_{(A)}^{X_{m}}(R^{\mathcal{X}}_{X_{m}})_{*}\nu\in{\mathcal{P}}({\mathcal{P}}({\Omega}_{(A)})^{X_{m}})$
if and only if there exists $\mu\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$ such that for any $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ and any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ we have
(5.11) $\widehat{ev}_{{\Omega}_{(A)}}^{(m)}\big(\nu^{X_{m}}_{(A)}\big)=P_{*}\pi_{(A)}^{X_{m}}(r^{\mathcal{X}}_{X_{m}})_{*}\mu\in{\mathcal{P}}({\Omega}_{(A)}^{X_{m}}).$
Equivalent, there exists $\nu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that (5.10) holds, if and only if for each $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ there exists $\mu_{X_{m}}\in{\mathcal{P}}({\mathcal{Y}}^{{\mathcal{X}}_{m}})$ such that for any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ we have
(5.12) $\widehat{ev}_{{\Omega}_{(A)}}^{(m)}\big(\nu^{X_{m}}_{(A)}\big)=P_{*}\pi_{(A)}^{X_{m}}(\mu_{X_{m}})\in{\mathcal{P}}({\Omega}_{(A)}^{X_{m}}),$
where the system $\{r_{X^{m}}^{X_{n}}:({\mathcal{Y}}^{X_{n}},\mu_{X_{n}})\to({\mathcal{Y}}^{X_{m}},\mu_{X_{m}}),\,X_{m}\leq X_{n}\in\mathrm{P_{fin}({\mathcal{X}})}\}$ of probability spaces is projective.
Remark 5.6.
(1) For the case $\mathcal{X}$} consists of one element, Theorem 5.5 is due to Orbanz [Orbanz2011, Theorem 1.1].
(2) Any $\mu\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$ can be written as
$\mu=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}(r^{\mathcal{X}}_{X_{m}})_{*}(\mu).$
By the Kolmogorov extension theorem, we can replace $(r^{\mathcal{X}}_{X_{m}})_{*}\mu$ in (5.11) by $\mu_{X_{m}}$ in (5.12) in the presence of the corresponding projective system, which is the content of the last “equivalence” assertion of Theorem 5.5.
Proof of Theorem 5.5.
(1) The equality (5.8) is a consequence of the functoriality $P_{*}:{\bf Probm}\to{\bf Meas}$ that assigns each measurable space $\mathcal{X}$} to measurable space ${\mathcal{P}}({\mathcal{X}})$ and each probabilistic morphism $T\in{\bf Probm}({\mathcal{X}},{\mathcal{Y}})$ to a measurable mapping $P_{*}T\in{\bf Meas}({\mathcal{P}}({\mathcal{X}}),{\mathcal{P}}({\mathcal{Y}}))$ [Giry82, Theorem 1].
The equality (5.9) follows from Lemma 5.4 and Remark 5.6 (2).
Now let us prove the last assertion of Theorem 5.5. The “only if” assertion is a consequence of (5.8) and (5.9).
Now we assume the “if” condition of Theorem 5.5. For each $X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$, we consider the projective system of probability spaces
[…]
together with mappings
[…]
Taking into account Orbanz’s result [Orbanz2011, Theorem 1.1] for ${\mathcal{P}}^{2}({\mathcal{Y}}^{X_{n}})$, see Remark 5.6(1), we conclude that there exists $\tilde{\nu}_{X_{n}}\in{\mathcal{P}}^{2}({\mathcal{Y}}^{X_{n}})$ such that
(5.13) $\mathfrak{m}^{n}(\nu^{X_{n}}_{(A)})=P_{*}^{2}\pi_{(A)}^{X_{n}}(\tilde{\nu}_{X_{n}}).$
Now we consider the following commutative system
(5.14) ${{\mathcal{P}}({\mathcal{Y}})^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{m}^{n}}$$\scriptstyle{(\pi_{(A)})_{*}^{X_{n}}}$${({\mathcal{P}}({\mathcal{Y}}^{X_{n}}),\tilde{\nu}_{X_{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(\pi_{(A)}^{X_{n}})}$${{\mathcal{P}}({\Omega}_{(A)})^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{m}^{n}}$${{\mathcal{P}}\big({\Omega}_{(A)}^{X_{n}},\mathfrak{m}^{n}(\nu^{X_{n}}_{(A)})\big).}$
Lemma 5.7.
The image $\mathfrak{m}^{n}({\mathcal{P}}({\mathcal{Y}})^{X_{n}})$ is a measurable subset of ${\mathcal{P}}({\mathcal{Y}}^{{\mathcal{X}}_{n}})$.
Proof of Lemma 5.7.
Let ${\mathcal{A}}_{\mathcal{Y}}$ be a countable algebra generating $\Sigma_{\mathcal{Y}}$. Then
$\mathfrak{m}^{n}({\mathcal{P}}({\mathcal{Y}})^{X_{n}})=\Big\{\mu\in{\mathcal{P}}({\mathcal{Y}}^{X_{n}}):\mu(A_{1}\times\ldots\times A_{n})=$
$\Pi_{i=1}^{n}\mu(\underbrace{{\mathcal{Y}}\times\ldots\times{\mathcal{Y}}}_{(i-1)\,times}\times A_{i}\times\underbrace{{\mathcal{Y}}\times\ldots\times{\mathcal{Y}}}_{(n-i)\,times})\text{ for any }A_{i}\in{\mathcal{A}}_{\mathcal{Y}},i\in\overline{1,n}\Big\}$
is a measurable subset of ${\mathcal{P}}({\mathcal{Y}}^{{\mathcal{X}}_{n}})$. This completes the proof of Lemma 5.7. ∎
Completion of the proof of Theorem 5.5. Since the map $\mathfrak{m}^{n}$ is injective, taking into account of Lemma 5.7, we conclude that
$\tilde{\nu}_{X_{n}}=(\mathfrak{m}^{n})_{*}\nu_{X_{n}}$
for some $\nu_{X_{n}}\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{X_{n}})$. Noting that the system $\{({\mathcal{P}}({\mathcal{Y}})^{X_{n}},\nu_{X_{n}}),X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})\}$ of probability spaces satisfies the condition of the Kolmogorov extension theorem, we complete the proof of Theorem 5.5. ∎
6. MacEachern’s Dependent Dirichlet Processes and Bayesian supervised learning
6.1. MacEachern’s Dependent Dirichlet Processes revisited
In this subsection, using Theorem 5.5, we revisit MacEachern’s Dependent Dirichlet Processes (DDPs) [MacEachern1999] [MacEachern2000] by synthesizing the categorical framework of this paper with the copula-based construction by Barrientos, Jara, and Quintana [BJQ2012]. As MacEachern [MacEachern2000] and Barrientos, Jara, and Quintana [BJQ2012], we assume that $\mathcal{Y}$} is a measurable subset of ${\mathbb{R}}^{n}$ and $\mathcal{X}$} is an arbitrary index set.
Let us first recall the general definition of a DDP from [BJQ2012, Definition 1]. A DDP is generated by a map (stochastic process)
$G:{\Omega}\times{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})$
where $({\Omega},P)$ is a probability space and for each $x\in{\mathcal{X}}$ the map $G(\cdot,x):({\Omega},P)\to{\mathcal{P}}({\mathcal{Y}})$ is measurable, or equivalently, the map
$\hat{G}:({\Omega},P)\to{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\hat{G}({\omega})(x):=G({\omega},x),$
is measurable. Furthermore, motivated by Sethuraman’s work [Sethuraman1994], $G$ must satisfy the following condition. For any $x\in{\mathcal{X}}$ and $B\in\Sigma_{\mathcal{Y}}$ we have
(6.1) $G({\omega},x)(B)=\sum_{i=1}^{\infty}W_{i}({\omega},x)\delta_{\theta_{i}({\omega},x)}(B),\,\text{ for $P$-a.e. }\;{\omega}\in{\Omega}$
where for all $x\in{\mathcal{X}}$ and $P$-a.e.   ${\omega}\in{\Omega}$
$W_{i}({\omega},x)=V_{i}({\omega},x)\prod_{j<i}(1-V_{j}({\omega},x)),$
with $V_{i}$ and $\theta_{i}:{\Omega}\times{\mathcal{X}}\to{\mathcal{Y}}$ described below.
  1. (1)
    $\{V_{i}:{\Omega}\times{\mathcal{X}}\to[0,1]\}_{i=1}^{\infty}$ where for each $x\in{\mathcal{X}}$ the sequence $\{V_{i}(\cdot,x):{\Omega}\to[0,1]\}_{i=1}^{\infty}$ are i.i.d. such that for any $i$
    $\big(V_{i}(\cdot,x)\big)_{*}P=\text{Beta}(1,\alpha(x))\in{\mathcal{P}}([0,1])\text{ where }\alpha(x)\in{\mathbb{R}}_{>0}.$
    The dependence structure of $V_{i}(\cdot,x)$ across $x\in{\mathcal{X}}$ for each $i$ is determined by a family ${\mathcal{C}}_{\mathcal{X}}^{V}:=\{C^{V}_{x_{1},\ldots,x_{d}}:[0,1]^{d}\to[0,1]\}$ of copula functions describing finite dimensional CDF of $(\hat{V}_{i})_{*}P\in{\mathcal{P}}([0,1]^{\mathcal{X}})$ where $\hat{V}_{i}:{\Omega}\to[0,1]^{\mathcal{X}},\hat{V}_{i}({\omega})(x):=V_{i}({\omega},x)$.
  2. (2)
    $\{\theta_{i}:{\Omega}\times{\mathcal{X}}\to{\mathcal{Y}}\}_{i=1}^{\infty}$ where for each $x$ the sequence $\{\theta_{i}(\cdot,x):{\Omega}\to{\mathcal{Y}}\}_{i=1}^{\infty}$ are i.i.d. such that for any $i$
    $\big(\theta_{i}(\cdot,x)\big)_{*}P=G^{0}_{x}\in{\mathcal{P}}({\mathcal{Y}}).$
    The dependence structure of $\theta_{i}(\cdot,x)$ across $x\in{\mathcal{X}}$ for each $i$ is determined by a family ${\mathcal{C}}_{\mathcal{X}}^{\theta}:=\{C^{\theta}_{x_{1},\ldots,x_{d}}:[0,1]^{d}\to[0,1]\}$ of copula functions describing finite dimenional CDF of $(\hat{\theta}_{i})_{*}P\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$ where $\hat{\theta}_{i}:{\Omega}\to{\mathcal{Y}}^{\mathcal{X}}$, $\hat{\theta}_{i}({\omega})(x):=\theta_{i}({\omega},x)$.
We denote the induced probability measure $(\hat{G})_{*}P\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ by ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}(\alpha_{\mathcal{X}}\in{\mathbb{R}}_{>0}^{\mathcal{X}},{\mathcal{C}}_{\mathcal{X}}^{\theta},{\mathcal{C}}_{\mathcal{X}}^{V},G^{0}_{\mathcal{X}}\in{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$. In fact, $(G,P)$ can be chosen as $\big(({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}(\alpha_{\mathcal{X}},\\ {\mathcal{C}}^{\theta}_{\mathcal{X}},C^{V}_{\mathcal{X}},G^{0}_{\mathcal{X}})\big)$ and $G$ is defined to be the natural evaluation mapping: $G({\omega},x)(B):={\omega}(x)(B)$ for any ${\omega}\in{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},x\in{\mathcal{X}}$ and $B\in\Sigma_{\mathcal{Y}}$.
According to Theorem 5.5, the probability measure ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}(\alpha_{\mathcal{X}},{\mathcal{C}}_{\mathcal{X}}^{\theta},{\mathcal{C}}_{\mathcal{X}}^{V},G^{0}_{\mathcal{X}})$ is uniquely determined by the projective system of its finite-dimensional projections. Let us describe this system. For any finite set of predictors $X_{m}=\{x_{1},\dots,x_{m}\}\subset{\mathcal{X}}$ and any finite measurable partition $(A)=(A_{1},\dots,A_{k})$ of $\mathcal{Y}$}, the corresponding projection is the probability measure
$\nu_{(A)}^{X_{m}}:=P^{2}_{*}\pi_{(A)}^{X_{m}}(R^{\mathcal{X}}_{X_{m}})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}(\alpha_{\mathcal{X}},{\mathcal{C}}_{\mathcal{X}}^{\theta},{\mathcal{C}}_{\mathcal{X}}^{V},G^{0}_{\mathcal{X}})\in{\mathcal{P}}({\mathcal{P}}({\Omega}_{k})^{X_{m}}).$
More explicitly, let
${\mathbf{p}}_{x_{i}}:=\big(G(\cdot,x_{i})(A_{1}),\ldots,G(\cdot,x_{i})(A_{k})):{\Omega}\to\Delta_{k}:={\mathcal{P}}({\Omega}_{k}).$
Then
$\nu_{(A)}^{X_{m}}=({\mathbf{p}}_{x_{1}},\ldots,{\mathbf{p}}_{x_{m}})_{*}P\in{\mathcal{P}}({\mathcal{P}}({\Omega}_{k})^{X_{m}}).$
The structure of this probability measure $\nu_{(A)}^{X_{m}}$ is as follows:
  • For any fixed $x_{i}\in X_{m}$, the marginal distribution $({\mathbf{p}}_{x_{i}})_{*}P\in{\mathcal{P}}^{2}({\Omega}_{k})$ is a Dirichlet distribution, as $\big(G(\cdot,x_{i})\big)_{*}P\in{\mathcal{P}}^{2}({\mathcal{Y}})$ is a Dirichlet process. Specifically,
    $({\mathbf{p}}_{x_{i}})_{*}P=\text{Dir}(\alpha(x_{i})G^{0}(x_{i})(A_{1}),\dots,\alpha(x_{i})G^{0}(x_{i})(A_{k})).$
  • The crucial point is that the joint distribution $\nu_{(A)}^{X_{m}}\in{\mathcal{P}}({\mathcal{P}}({\Omega}_{k})^{X_{m}})$ is not a simple product of these marginal Dirichlet distributions $({\mathbf{p}}_{x_{i}})_{*}P\in{\mathcal{P}}^{2}({\Omega}_{k})$. The dependence between ${\mathbf{p}}_{x_{i}}$ and ${\mathbf{p}}_{x_{j}}$ for $i\neq j$ is induced by the dependence structure of the underlying stick-breaking processes $\{V_{l}(x)\}_{l=1}^{\infty}$ and $\{\theta_{l}(x)\}_{l=1}^{\infty}$. This dependence is precisely what is encoded by the copula families ${\mathcal{C}}_{\mathcal{X}}^{V}$ and ${\mathcal{C}}_{\mathcal{X}}^{\theta}$.
Furthermore, Theorem 5.5 provides a consistency condition involving a “mean” measure $\mu\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$. For the ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$, this corresponds to the map of base measures $G^{0}_{\mathcal{X}}:{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})$ defined by $x\mapsto G^{0}_{x}$, regarded as an element in ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\stackrel{{\scriptstyle\mathfrak{m}^{\mathcal{X}}}}{{\hookrightarrow}}{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$. The ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ is centered around this collection of measures, as ${\mathbb{E}}_{P}[G_{x}]=G^{0}_{x}$. The projective system $\{\nu_{(A)}^{X_{m}},X_{m}\in(\mathrm{P_{fin}}({\mathcal{X}}),\leq)\}$ must satisfy the condition (5.11):
$\widehat{ev}_{{\Omega}_{k}}^{(m)}(\nu_{(A)}^{X_{m}})=P_{*}\pi^{X_{m}}_{(A)}\circ(r^{\mathcal{X}}_{X_{m}})_{*}(\mathfrak{m}^{\mathcal{X}}G^{0}_{\mathcal{X}}).$
This simply states that the expected value of the random vector ${\mathbf{p}}_{x_{i}}$ is the vector of probabilities of the base measure, $(G^{0}_{x_{i}}(A_{1}),\dots,G^{0}_{x_{i}}(A_{k}))$, which is a fundamental property of the Dirichlet process [GV2017, §4.1.4].
In summary, we can characterize MacEachern’s ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ in the following way:
Theorem 6.1.
The law of a Dependent Dirichlet Process, ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}(\alpha_{\mathcal{X}},{\mathcal{C}}_{\mathcal{X}}^{\theta},{\mathcal{C}}_{\mathcal{X}}^{V},G^{0}_{\mathcal{X}})$, is the unique probability measure $\nu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ that satisfies the two conditions of Theorem 5.5, where:
  1. (1)
    The projective system of second-order probability measures $\{\nu_{(A)}^{X_{m}}\}$ is defined such that each $\nu_{(A)}^{X_{m}}$ is the law of an $m$-tuple of measurable mappings $\{{\mathbf{p}}_{x_{i}}:{\Omega}\to{\mathcal{P}}({\Omega}_{k})\}_{i=1}^{m}$, where the marginal law $({\mathbf{p}}_{x_{i}})_{*}P\in{\mathcal{P}}^{2}({\Omega}_{k})$ is Dirichlet, and their joint dependence structure is determined by the copula families ${\mathcal{C}}_{\mathcal{X}}^{V}$ and ${\mathcal{C}}_{\mathcal{X}}^{\theta}$.
  2. (2)
    The corresponding projective system of first-order measures is given by the base measure map $\mu=\mathfrak{m}^{\mathcal{X}}\circ G^{0}_{\mathcal{X}}\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$.
The same reasoning can be applied to characterize the simpler “single-weights” and “single-atoms” ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$s from [BJQ2012, Definitions 2, 3] by appropriately simplifying the copula structure (e.g., using independence copulas for the weights in the single-weights case).
6.2. Posterior predictive distributions of Bayesian supervised learning models $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}},\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$
Let us consider a Bayesian supervised learning model $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}},\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},\\ {\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ where ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}={\mathcal{D}}{\mathcal{D}}{\mathcal{P}}(\alpha_{\mathcal{X}},{\mathcal{C}}^{\theta}_{\mathcal{X}},{\mathcal{C}}^{V}_{\mathcal{X}},G^{0}_{\mathcal{X}})$ is described in Subsection 6.1. In particular, $\mathcal{Y}$} is a measurable subset in ${\mathbb{R}}^{k}$. To compute the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$ we restrict ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ to ${\mathcal{P}}({\mathcal{Y}})^{A}$ where $A=[T_{m}]\cup[\pi_{\mathcal{X}}(S_{n})]$. The restriction $({\mathbb{R}}^{\mathcal{X}}_{A})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ of ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ to ${\mathcal{P}}({\mathcal{Y}})^{A}$ is ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}(\alpha_{A},{\mathcal{C}}^{\theta}_{A},{\mathcal{C}}^{V}_{A},G^{0}_{A})$ where
$\alpha_{A}=(\alpha_{\mathcal{X}})_{|A},{\mathcal{C}}^{\theta}_{A}=({\mathcal{C}}^{\theta}_{\mathcal{X}})_{|A},{\mathcal{C}}^{V}_{A}=({\mathcal{C}}^{V}_{\mathcal{X}})_{|A},G^{0}_{A}=(G^{0}_{\mathcal{X}})_{|A}.$
Next, we shall apply Theorem 4.1 to compute ${\mathcal{P}}_{T_{m}|S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$ where $T_{m}=(t_{1},\ldots,t_{m})$ and $\Pi_{\mathcal{X}}(S_{n})=(x_{1},\ldots,x_{n})$. We have
(6.2) $\mu^{0}_{T_{m},S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}=\int_{{\mathcal{P}}({\mathcal{Y}})^{A}}\otimes_{i=1}^{m}{\mathcal{D}}{\mathcal{P}}(h(t_{i})))\otimes_{j=1}^{n}{\mathcal{D}}{\mathcal{P}}(h(x_{j}))d(R^{\mathcal{X}}_{A})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}(h)$
where ${\mathcal{D}}{\mathcal{P}}(h(t_{i}))\in{\mathcal{P}}({\mathcal{Y}})$ is the marginal Dirichlet process of ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ evaluated at $t_{i}$. Note that $\mu^{0}_{T_{m},S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}\in{\mathcal{P}}({\mathcal{Y}})^{m+n}\subset{\mathcal{P}}({\mathbb{R}}^{k})^{m+n}$ has the marginal probability measures defined by ${\mathcal{D}}{\mathcal{P}}(h(t_{i}))$ and ${\mathcal{D}}{\mathcal{P}}(h(x_{j}))$. Thus to determine the joint measure $\mu^{0}_{T_{m},S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$ it suffices to compute its copula functions $C_{T_{m},S_{n},({\mathbb{R}}^{\mathcal{X}}_{A})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$. Knowing $C_{T_{m},S_{n},(R^{\mathcal{X}}_{A})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$, we can apply Theorem 3.4.1 in [DS2016] to compute ${\mathcal{P}}_{T_{m}|S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$, using the recursive formula in Theorem 4.1.
Alternatively, to compute the posterior predictive distribution ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}_{T_{m}|S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$, we may apply Theorem 5.5 and Theorem 3.4, or the following Theorem.
Theorem 6.2.
Let $\mathcal{Y}$} be a measurable space, ${\mathcal{X}}:=\{x_{1},\ldots,x_{n}\}$ a finite set, $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, and $X_{n}=\Pi_{{\mathcal{X}}}(S_{n})\in{\mathcal{X}}^{n}$, $T_{m}\in{\mathcal{X}}^{m}$. Assume that for any $(A)\in\pi({\mathcal{Y}})$ there exists a Markov kernel ${\mathbf{q}}^{n}_{(A),m}:{\Omega}_{(A)}^{n}\to{\mathcal{P}}({\Omega}_{(A)}^{m})$ which is a regular conditional probability measure of the joint distribution of ${\mathbf{p}}^{(A)}_{T_{m},X_{n}}(P_{*}(\pi_{(A)})_{*}^{\mathcal{X}}\mu)$ such that the following diagram is commutative for any $(A)\leq(B)\in\pi({\mathcal{Y}})$.
${{\mathcal{P}}({\Omega}_{(B)}^{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(\pi^{(B)}_{(A)})^{m}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(B)}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}_{(B),m}^{n}}$$\scriptstyle{(\pi^{(B)}_{(A)})^{n}}$${{\mathcal{P}}({\Omega}_{(A)}^{m})}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}^{n}.}$$\scriptstyle{{\mathbf{q}}_{(A),m}^{n}}$
Assume that there exists a Markov kernel ${\mathbf{q}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ such that for any $(A)\in\pi({\mathcal{Y}})$ the following diagram is commutative
${{\mathcal{P}}({\mathcal{Y}}^{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(\pi_{(A)})^{m}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\mathcal{Y}}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}^{n}_{m}}$$\scriptstyle{(\pi_{(A)})^{n}}$${{\mathcal{P}}({\Omega}_{(A)}^{m})}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}^{n}.}$$\scriptstyle{{\mathbf{q}}^{n}_{(A),m}}$
Then ${\mathbf{q}}^{n}_{m}(\Pi_{\mathcal{Y}}(S_{n}))={\mathcal{P}}_{T_{m}|S_{n},\mu}$.
This Theorem is proved in the same way as Theorem 3.4, so we omit its proof. s
7. Final remarks
  1. (1)
    In this paper we proved that batch Bayesian learning equals Bayesian online learning under the assumption of conditionally independent data, making Bayesian learning more efficient in the presence of complex data. Our Theorem 4.1 generalizes the celebrated Kalman filter.
  2. (2)
    Bayesian regression learning with corrupted measurements can be extended to nonlinear spaces $\mathcal{Y}$} where we can model measurement error using probability measures, e.g., for homogeneous Riemannian manifolds $\mathcal{Y}$}. Corollary 4.2 can be extended for Bayesian regression learning with corrupted measurements in the same way.
  3. (3)
    It is important to find a suitable concept of predictive consistency of Bayesian supervised learning which would agrees with the classical concept and the concept in a recent work by P. Koerpernik and F. Pfaff. A possible solution is to introduce the notion of predictive consistency at a finite subset $A\subset{\mathcal{X}}$ and consider posterior predictive distributions of the form ${\mathcal{P}}_{T_{m}|S_{n},\mu}$ where $T_{m}]\subset A$ and $\Pi_{\mathcal{X}}(S_{n})]\in A$, moreover, $\Pi_{\mathcal{X}}(S_{n})$ visits each element of $A$ infinitely many times. If ${\mathcal{X}}=1$ and the sampling operator is Markov kernel this concept is the notion of posterior consistency in classical Bayesian statistics [GV2017, §6.8.3]. Furthermore, this notion also agrees with the condition of recurrent density entered in the concept of posterior consistency of Gaussian process regressions in [KP2021].
Acknowledgement
Research of HVL was supported by the Institute of Mathematics, Czech Academy of Sciences (RVO: 67985840). The author would like to thank Steven MacEachern for suggesting the paper by Barrientos-Jara-Quintana to her. She also thanks Xuan Long Nguyen for inviting her to the workshop “Bayesian Modeling, Computation, and Applications” at Ho Chi Minh City in July 2025, which stimulated her working over this paper.